### v1

# Accelerating Mobile Language Model Generation via Hybrid Context and Hardware Coordination

Abstract.
Enhancing on-device large language models (LLMs) with contextual information from local data enables personalized and task-aware generation, powering use cases such as intelligent assistants and UI agents. While recent developments in neural processors have substantially improved the efficiency of prefill on mobile devices, the token-by-token generation process still suffers from high latency and limited hardware utilization due to its inherently memory-bound characteristics. This work presents CoordGen, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices. The framework introduces three synergistic components: (1) adaptive execution scheduling, which dynamically balances compute graphs between prefill and decoding phases; (2) context-aligned drafting, which improves speculative efficiency through lightweight online calibration to current tasks; and (3) hardware-efficient draft extension, which reuses and expands intermediate sequences to improve processing parallelism and reduce verification cost. Experiments on multiple smartphones and representative workloads show consistent improvements of up to 3.8× in generation speed and 4.7× in energy efficiency compared with existing mobile inference solutions. Component-level analysis further validates the contribution of each optimization.
1. Introduction
Recent advances in large language models (LLMs) (Brown et al., 2020; OpenAI et al., 2024; Touvron et al., 2023; Guo et al., 2024), rising concerns over privacy, and rapid progress in hardware have fueled interest in on-device LLM. In particular, the integration of neural processing units (NPUs) (Qualcomm, 2024, 2025b) makes it feasible to run LLMs directly on mobile devices (Google, 2024; Huawei, 2024; Samsung, 2024; Apple, 2024a), enabling applications such as voice assistants and UI automation (Apple, 2024b; AI, 2024; Wen et al., 2024; Lee et al., 2024). To extend LLMs’ capabilities in handling complex tasks, context-augmented generation (CAG) has emerged as a key technique (Wang and Chau, 2024; Ouyang et al., 2025). As shown in Figure 1, CAG augments a user’s query with task-relevant contexts before inference to strengthen the knowledge grounding of LLM. With CAG, LLMs can, for example, summarize documents using user-uploaded files, generate responses by retrieving and integrating information from the chat history, and act as UI agents by reasoning over current UI screenshots captured from the operating system, significantly expanding their functionality.
Despite these benefits, on-device CAG still has high end-to-end latency, primarily from decoding (accounting for 73.5% of total inference, details in Section 2.1). When diving into the reasons, we find that the prefill bottleneck is largely eliminated by NPU-based chunked prefill (Xu et al., 2025b; Chen et al., 2025a; Xue et al., 2024) (only 2 s for 1–2K tokens context). In contrast, the decoding stage remains memory-bound, a scenario poorly suited for NPUs. Therefore, most existing work resorts to CPU execution, leading to high latency. Motivated by the NPU’s superior energy efficiency and lower contention, this paper raises the question whether NPUs can be effectively utilized to accelerate the decoding stage.
Figure 1. Workflow of context-augmented generation and comparison between CoordGen and related work :
To address this issue, we observe that CAG outputs exhibit strong similarity to both the interaction history and the augmented context (Section 2.2), making them well-suited for retrieval-based speculative decoding (R-SD), a technique originally developed for code generation, and in turn, to NPU acceleration. In many CAG tasks, generated content is semantically close to previously produced tokens and to the augmented context: for example, summarization and automated email reply often reuse spans from user-uploaded documents or earlier emails, while UI agents repeatedly emit fixed commands (e.g., “click set clock button”). These characteristics align naturally with R-SD, which proceeds in two stages: (1) drafting, where a retriever proposes candidate token sequences by selecting similar fragments from the augmented context or a local store (e.g., a code repository); and (2) verifying, where the LLM compares these drafts against its own next-token predictions and corrects discrepancies (see Section 2.2). By evaluating multiple drafts in parallel, R-SD shifts decoding from a memory-bound to a compute-bound workload, making it a strong fit for NPU acceleration.
In this paper, we present CoordGen, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading. During prefill, CoordGen employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification. Designing CoordGen requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
$\bullet$ NPU execution: costly switching overhead for distinct prefill and decoding optimal compute graphs. Prefill favors large-shape graphs with long sequence inputs which degrade decoding performance, while decoding benefits from small-shape graphs that underperform for prefill. Since the NPU only supports inference on static shapes, achieving optimal end-to-end performance thus requires switching between prefill-optimal and decoding-optimal compute graphs, incurring a significant performance overhead. For instance, our preliminary evaluation on Qwen2.5-1.5B model shows this switching overhead incurs a 94.9% delay of end-to-end latency.
$\bullet$ Draft construction: task context and LLM outputs follow different lexical distributions, yielding a low acceptance ratio. In R-SD, the LLM only accepts tokens that exactly match its own prediction. Our preliminary evaluation shows that drafts constructed from the task context often have high semantic similarity to the target output yet exhibit notable lexical gaps (e.g., normalized Levenshtein distance >0.11>0.11 for Qwen2.5-1.5B). As a result, LLM often accepts only few tokens per draft, falling back to token-by-token decoding and underscoring the need for lexical distribution calibration.
$\bullet$ Draft verification: short or absent drafts fail to saturate NPU computation. Mobile NPUs excel at processing large batches or long sequences due to their weight-stationary design where short inputs obtain fewer benefits (Jouppi et al., 2017; Chen et al., 2025a). However, due to strict suffix-matching, existing R-SD methods often generate only a few candidate tokens or none at all. Such short sequences cannot fully exploit the NPU’s weight-stationary design, resulting in low verification throughput. For example, our experiments show that over 75% of the drafts fall below length 8, which is far below the verification-optimal length (64) and leads to 50% lower throughput.
To address these challenges, we propose three techniques:
$\bullet$ Progressive Graph Scheduling (Section 3.2). CoordGen deploys two specialized graphs and switches between them asynchronously to minimize overhead. The LLM is partitioned into blocks. During each round of chunked prefill, we switch to a decoding-optimized block graph while prefill computation continues in parallel. The switched decoding graph is invoked multiple times within the prefill stage to ensure result correctness. This design rests on two observations: (i) executing short-input graphs multiple times is equivalent to a single long-input execution, and (ii) graph loading can be overlapped with one-round chunked prefill computation. Consequently, CoordGen can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
$\bullet$ In-Context Distribution Calibration (Section 3.3). Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, CoordGen uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search. In the subsequent retrieval-based drafting stage, CoordGen retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations. Compared with costly precomputation or task-specific fine-tuning, this lightweight procedure incurs negligible overhead.
$\bullet$ NPU-Optimized Draft Reuse (Section 3.4). CoordGen extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy. This is motivated by the observation that rejected tokens may still capture correct semantics and be accepted in later decoding steps. By reusing such tokens, CoordGen increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
[…]
Our contributions are summarized as follows:
$\bullet$ We present the first study of inefficiencies in context-aware generation for mobile tasks, identifying key optimization opportunities of using retrieval-based speculative decoding.
$\bullet$ We design and implement CoordGen, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that CoordGen delivers consistent and significant performance gains in diverse mobile scenarios.
2. Background
2.1. On-device Context-Augmented Generation for Mobile Task
Context-augmented generation. With the rapid development of LLMs and increasing demands for privacy protection, there is a growing trend of deploying large models on mobile devices (Belcak et al., 2025; Navardi et al., 2025; Union, 2021; Kshetri, 2023; Eleftherakis et al., 2024). An emerging paradigm, context-augmented generation, leverages both the long-context capability of LLMs and the abundance of private on-device data, showing strong potential for supporting complex tasks such as personal assistant and UI automation (Wang and Chau, 2024; Ouyang et al., 2025; Park et al., 2025). Typically, context-augmented generation first exploits additional information, either directly provided by the user or retrieved from relevant sources, to enrich the prompt with task-specific contexts. Then, the normal LLM inference procedure is executed, including prefill and decoding stages. The enriched prompt can alleviate the limitations of the model’s built-in knowledge and enable it to handle a wider range of concrete tasks. For instance, in a summarization task, the user may supply a meeting transcript as context, and the LLM directly generates a concise summary from it. In another case, a personalized assistant retrieves a user’s past chat records to draft a reply in the user’s typical style, where the retrieved records provide the key context for generation.
LLMs on NPUs. As modern mobile SoCs increasingly integrate high-performance NPUs (Table 1), recent mobile LLM frameworks aim to offload LLM inference onto NPUs to improve efficiency. In practice, NPUs are mainly used for the prefill stage, which benefits from their high throughput on long input sequences, while decoding is still executed on CPUs or GPUs (Xu et al., 2025b; Chen et al., 2025a; Xue et al., 2024). Our experiments reveal that mobile CAG still suffers from high latency under this design, with decoding dominating end-to-end performance. For example, for Qwen2.5-1.5B (Qwen et al., 2025) running on a OnePlus 13, decoding accounts for 90.2% of latency in question answering (QA) tasks and 87.2% in summarization, even when prefill runs at over 1000 tokens/s (Xu et al., 2025b) (Figure 2). Simply offloading decoding to the NPU provides no advantage over CPU execution and often increases overall latency, as decoding is inherently memory-bound, processing one token at a time and leaving the NPU underutilized.
Figure 2. Comparison of prefill and decoding latency for several CAG tasks. Task details are provided in Section 4.1.
Table 1. Specifications of representative mobile NPUs.
Vendor SoC NPU Perf. (Tops)
Qualcomm 8 Gen 3 Hexagon NPU 73
Apple A18 Neural Engine 35
MediaTek K9300 APU 790 60
Huawei Kirin-9000 Ascend NPU 16
Perf. = INT8 Performance in Tops.
These findings are closely tied to the architectural characteristics of mobile NPUs. First, their systolic-array with weight-stationary design excels at long input sequences: weights are preloaded once and reused across tokens, greatly reducing memory traffic. This property aligns naturally with the prefill stage, which requires lots of computation over long prompts. Second, NPUs rely on static compute graphs with precompiled execution plans, where operator arrangements and tensor shapes are fixed in advance for efficiency. Naively reusing prefill-optimized graphs for decoding leads to severe inefficiencies. As a result, while NPUs eliminate prefill bottlenecks, they bring little benefit to decoding, which thus remains the dominant source of latency and energy consumption in on-device CAG.
2.2. Opportunity: Retrieval-based Speculative Decoding
A key observation is that CAG naturally inherits similarity from the request or the augmented context, making it well-suited for retrieval-based speculative decoding (Saxena, 2023; He et al., 2024; Hu et al., 2024; Somasundaram et al., 2024), an acceleration technique exploiting similarity between contexts and outputs. We find that mobile CAG tasks exhibit the same property, offering strong opportunities for acceleration via R-SD.
$\bullet$ History similarity in similar tasks: outputs from similar tasks resemble previous generations because their contexts are similar. For example, a UI automation agent often repeats commands like clicking the same button. Our evaluation on four UI automation task groups (Wen et al., 2024) with Qwen2.5-1.5B (Qwen et al., 2025) confirms this (Figure 3(a)): predictions within the same task show the highest semantic similarity (0.72), those within the same group moderate similarity (0.61), and across groups low similarity (0.38). Cosine similarity is computed using embeddings from all-MiniLM-L6-v2 (Wang et al., 2020).
$\bullet$ Context similarity at semantic level: Outputs align with the enriched context, which often includes direct clues for response (Wang and Chau, 2024; Ouyang et al., 2025; Park et al., 2025). For instance, assistants answering QA queries frequently extract information from documents. On a summarization dataset (Xia et al., 2024) with Qwen2.5-1.5B, we compute semantic similarity between context and predictions (via cosine similarity) and lexical overlap (via Levenshtein distance). Figure 3(b) shows that 90% of samples exhibit high semantic similarity (>0.76>0.76) but low lexical overlap (>0.11>0.11), reflecting a stylistic gap between context and outputs.
(a) History
(b) Context
Figure 3. Demonstration of similarities in context-augmented generation.
Figure 4. Workflow of retrieval-based speculative decoding.
R-SD exploits these similarities through two stages (Figure 4): (1) retrieval-based drafting, where relevant fragments are retrieved from history or context to draft candidate sequences. For example, tokens “need to” may retrieve drafts such as “click on” and “find the”; (2) verifying, where the LLM validates these candidates in parallel, accepting matches and discarding mismatches to guarantee correctness (Leviathan et al., 2023a). In the example where model expects “find the”, mismatched drafts like “click on” are discarded while the matched draft is accepted, extending the output to “find the Google”. The similarities in mobile CAG tasks offer promising draft candidates, enabling R-SD to efficiently generate multiple tokens in parallel. R-SD thus substantially raises the computation-to-memory ratio, shifting decoding from memory-bound to compute-bound where NPU excels.
Parallel to retrieval-based methods, model-based speculative decoding (M-SD) (e.g., EAGLE (Li et al., 2024c, b), Medusa (Cai et al., 2024)) trains auxiliary draft models to generate candidates, incurring extra training, storage, and inference overheads. R-SD avoids retraining by directly using contextual information of high similarity, achieving draft latencies under 10ms (Hu et al., 2024). Therefore, R-SD is more suitable for resource-constrained devices while M-SD suffers from higher latency, as shown in our evaluations (Section 4.2).
2.3. Challenges: Aligning Retrieval-based Speculative Decoding with NPU
Although R-SD offers a promising approach to accelerate CAG with NPU offloading, integrating it with NPU offloading poses several challenges.
Figure 5. Trade-off between prefill and decoding with fixed-size compute graphs, highlighting the importance to deploy specialized graphs for each stage.
$\bullet$ NPU execution: static compute graph design hinders performance improvement for both prefill and decoding. Static compute graph constraints fixed input and output tensor shapes during execution. To support variable-length sequences in LLM inference, existing systems adopt a chunking strategy, i.e., splitting inputs into fixed-size segments (Xu et al., 2025b; Chen et al., 2025a). However, a unified chunk size for both prefill and decoding yields suboptimal performance. Figure 5 compares prefill and speculative decoding speed using different compute graphs on a Redmi K70 Pro smartphone, with a fixed acceptance ratio of 4.2 (based on evaluation results, see Section 4.5). Results show that compute graphs with larger chunk size improve prefill throughput but hurt decoding performance. For instance, despite a fast prefill speed (¡8 ms/token) of Qwen2.5 1.5B using graph with chunk size 32, the decoding speed using the same graph drops to about 160 ms/token. This tradeoff necessitates specialized compute graphs to improve performance for both prefill and decoding. However, on resource-constrained devices, naively maintaining multiple graphs in memory is costly, highlighting the needs of lightweight graph scheduling algorithms.
Figure 6. Comparison of verification speed with different draft lengths, and the length distribution of drafts generated by existing R-SD method (Saxena, 2023) (histogram).
$\bullet$ Draft construction: divergence between task context and LLM output reduces the efficiency of speculative decoding. As discussed in Section 2.1, contexts and predictions often exhibit lexical gaps (i.e., normalized Levenshtein distance >0.11>0.11). These gaps result in frequent rejection of retrieved drafts, despite their success in capturing the similar semantics as LLM. This degrades decoding to a token-by-token process where NPUs offer little speedup due to underutilization. Lexical gaps arise from divergence between the task context and the LLM’s distribution, necessitating alignment of the two distributions.
$\bullet$ Draft verification: existing R-SD methods underutilize NPUs’ weight-stationary design due to drafting short sequences. The design of mobile NPUs, while ideal for compute-intensive operations such as prefill, is less effective when input sequences become short in the decoding stage. Existing R-SD methods perform retrieval by strict suffix matching between context and output (Somasundaram et al., 2024; He et al., 2024; Hu et al., 2024), which frequently results in short or empty drafts. Figure 6 demonstrates the influence of draft length on verification speed and the length distribution of generated draft on a summarization dataset (Xia et al., 2024) . Results show that over 90% of drafts fall below the NPU-unfriendly length, significantly reducing the verification speed as NPU computing capacity is underutilized. This results in a similar speed to CPU (shown in the horizontal line). For instance, the verification speed of Qwen2.5 1.5B drops from 5.7 ms/token to 51.1 ms/token when the draft length decreases from 32 to 1.
3. Methodology
3.1. Overview
Figure 7. Overview of CoordGen framework.:
Design goal. CoordGen accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding. This entails (1) deploying and scheduling two compute graphs optimized for prefill and decoding, (2) aligning task context with LLM distribution, and (3) enriching drafts with sufficient high-value tokens to fully utilize NPU resources.
Workflow. Figure 7 illustrates the system architecture and workflow of CoordGen.
  • Offline stage. CoordGen partitions the LLM into multiple blocks, each containing several transformer layers. It profiles the load and compute latency of each block through warm-up to determine scheduling strategy.
  • Online stage. Upon receiving a request, CoordGen fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph. During prefill, progressive graph scheduling (Section 3.2) incrementally switches the prefill-optimized graph to the decoding-optimized graph with load overlapped by computation. After prefill, output logits are used for in-context distribution calibration (Section 3.3), augmenting the database with calibrated contexts. CoordGen then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database. It iteratively constructs a drafting tree and verifies it via a confidence-based strategy to identify and reuse valuable tokens (Section 3.4). After an end-of-sequence token is predicted, the final response is returned and stored in the database manager as a historical context for future requests.
Key techniques. CoordGen incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
  • Progressive NPU graph scheduling (Section 3.2) enables asynchronous graph switching by overlapping load and computation across chunked prefill rounds.
  • In-context distribution calibration (Section 3.3) aligns task context with model expectations using prefill logits.
  • NPU-optimized draft reuse (Section 3.4) maximizes NPU utilization and reduces verification costs by reusing plausible tokens from rejected drafts.
3.2. Progressive Graph Scheduling
Challenges of fixed-shape compute graphs. As discussed in Section 2.3, mobile NPUs require fixed-shape compute graphs, forcing the prefill input to be chunked for variable-length requests. This constraint also forces prefill and decoding to share the same graph despite their distinct input patterns: prefill processes long sequences (e.g., chunk size 256), while decoding handles short inputs (e.g., chunk size 32). Using a unified graph thus leads to inefficiency. To address this, we consider two specialized graphs: $G^{1}$ optimized for prefill and $G^{2}$ optimized for decoding. We denote the computation latency of a forward pass of model $\operatorname{P}$ on graph $G^{x}$ ($x\in\{1,2\}$) as $T_{\text{compute}}(\operatorname{P};G^{x})$.
[…]
Switching graph progressively. CoordGen gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
  • Graph equivalence by repetition: executing $G^{2}$ multiple times on shorter inputs yields the same result as one execution of $G^{1}$ on longer inputs, allowing early switching. As CoordGen maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
  • Chunk-level scheduling: multiple block computation can be overlapped with a single block loading.
(a) Progressive switching overlaps $G^{2}$ loading with computation across blocks.
[…]
Figure 8. Progressive graph scheduling example with $N=6$ and two chunks. (a) shows the workflow of chunk 1 computation; (b) compares prefill pipelines under different scheduling schemes.
Figure 8 illustrates progressive scheduling with $N=6$ and two input chunks. We suppose that CoordGen decides to overlap loading a block by two blocks of computation based on the profiling results. Initially, all blocks are loaded with $G^{1}$. CoordGen computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$. Then $\operatorname{P}_{3},\operatorname{P}_{4}$ execute while $G^{2}$ loads for $\operatorname{P}_{5}$. To compute $\operatorname{P}_{5},\operatorname{P}_{6}$ under $G^{2}$, chunk 1 is split into sub-chunks matching $G^{2}$’s shape, with $G^{2}$ executed repeatedly. Prefill of chunk 1 finishes with output logits while half of the model is switched. The same procedure applies to chunk 2. In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency (Figure 8(b)).
Problem Formulation. Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, CoordGen aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round. The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
Scheduling methodology. This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (Korf, 1998) (proofed in Appendix A). CoordGen addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute. This strategy enables fully masking the load overhead by computation to minimize switching overhead. CoordGen finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results. To hide load latency, the sum of the next $k$ blocks’ compute time must be larger than the load overhead of the last block not switched, formulated as:
(1) $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})\;\leq\;\sum_{\ell=0}^{k-1}\widehat{T}_{\text{compute}}(\operatorname{P}_{i+\ell};G^{x}_{i+\ell}).$
CoordGen greedily chooses the minimal feasible $k$ that satisfies Equation 1:
(2) $k^{\star}\;=\;\min\Big\{k\geq 1\;\big|\;\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})\leq\sum_{\ell=0}^{k-1}\widehat{T}_{\text{compute}}(\operatorname{P}_{i+\ell};G^{x}_{i+\ell})\Big\}.$
Then CoordGen executes $\operatorname{P}_{i},\dots,\operatorname{P}_{i+k^{\star}-1}$ under current graphs while loading $G^{2}$ for $\operatorname{P}_{j}$. This process repeats until all blocks are switched.
3.3. In-Context Distribution Calibration
Challenges of language divergence. As discussed in Section 2.1, leveraging similarity on mobile LLMs is hindered by distribution divergence between task context and LLM modeling. Models often interpret the context with different lexical patterns, thus reducing the acceptance ratio of drafts directly generated from the context. These lexical gaps appear prominently at the token level. For example, in Figure 9, a user request states “… Clinton’s security detail has arrived at Des Moines …,” while the LLM, tasked with summarizing, prefers “Clinton arrived at Des Moines.” For classical retrieve-based speculative decoding approaches (Saxena, 2023; He et al., 2024; Hu et al., 2024), draft “Clinton’s security” is rejected due to lexical mismatch between token ’s and target arrived. This results in only one token is generated (arrived) by one forward pass as the following draft security is discarded, reducing decoding efficiency. Such mismatch problem arises in various mobile scenarios such as auto-reply or RAG-based QA, where human-written contexts diverge from the model’s distribution.
Figure 9. Overview of in-context distribution calibration with a summarization example. :
To tackle this issue, one may forcibly accept drafts with lexical gaps based on semantic-level similarity. However, setting a stringent similarity threshold results in a low acceptance ratio with only marginal performance gains, whereas a loose threshold causes non-negligible accuracy loss, e.g., drops from 0.9 to 0.75 if candidates are forcibly accepted ignoring the distribution divergence (Holsman et al., 2025; Bachmann et al., 2025). This necessitates calibration of draft and model distributions. A straightforward solution is to precompute a calibrated database by running large volumes of user inputs through the LLM, which suffers from poor generalization to unseen queries and requires prohibitive storage and offline processing. Another method is to fine-tune the LLM to adapt to specific task scenarios, which also consumes lots of resources and energy and is hard to cover the wide diversity of mobile tasks.
Leveraging prefill logits. To avoid heavy offline or fine-tuning costs, our key observation is that after prefill completes, next-token logits for each token in the entire context have already been calculated. These logits reflect how LLM encodes task semantics and can be used to approximate the model distribution. To that end, we can build a calibrated database for drafting with negligible performance overhead. Therefore, we propose in-context distribution calibration, an online, lightweight method that dynamically aligns context with the LLM distribution using prefill logits.
Workflow. After prefill, CoordGen first collects predicted logits for all positions. Then CoordGen builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
[…]
Overhead analysis. The main costs come from sampling and storing calibrated results. In our experiments, sampling incurs an amortized latency of about 2 ms per input token, while building and maintaining calibrated tokens peak below 500 MB memory for a 3B model (Section 4.3).
3.4. NPU-Optimized Draft Reuse
Challenges of short potentially acceptable drafts. Existing studies on R-SD focus on improving draft quality, proposing techniques such as pruning draft tree by estimating token confidence (Somasundaram et al., 2024; Hu et al., 2024; Wu et al., 2025a). However, as discussed in Section 2.3, a large proportion of drafts generated by existing methods have short lengths, which is harmful to the NPU efficiency. Naively lengthening drafts by loosing retrieving conditions cannot obtain performance benefits as irrelevant tokens will be rejected in verification.
Reuse plausible drafts. As highlighted in Section 2.1, the model output usually shares similar semantics with the task context despite existing lexical gaps. Although these gaps cause rejection due to token-level misalignment, the rejected tokens include correct semantics and may be accepted in the next verification. CoordGen identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
Table 2. Given draft $T,x_{1}\dots x_{n}$ which misaligns with target $z_{1}\dots z_{n+1}$ at position $e$, the possible patterns of segment $z_{e}\dots z_{n}$ and corresponding reusable tokens. In cases missing and synonymous, tokens after $x_{e+\epsilon}$ are omitted for simplicity.
Case Pattern Reusable Ratio
Missing $z_{e}\dots z_{e+\delta}x_{e}\dots x_{e+\epsilon}$ $x_{e}\dots x_{e+\epsilon}$ 15.4%
Synonym $z_{e}\dots z_{e+\delta}x_{e+\gamma}\dots x_{e+\epsilon}$ $x_{e+\gamma}\dots x_{e+\epsilon}$ 23.1%
Redundant $z_{e}\dots z_{n}$ $\emptyset$ 61.5%
Rejection causes analysis. We first analyze the rejection reasons to explore how to identify plausible tokens. Let $\mathbf{x}=[T,x_{1},\dots,x_{n}]$ be the draft sequence where $T$ is the most recently accepted token, and $\mathbf{y}=[y_{1},\dots,y_{n+1}]$ be the model’s predictions. Suppose the expected sequence is $\mathbf{z}=[z_{1},\dots,z_{n+1}]$ and rejection occurs at position $e$ due to a mismatch: $x_{e}\neq y_{e}$. Thus, the accepted sequence is $x_{1},\dots,x_{e-1},y_{e}$. According to the patterns of the expected sequence $z_{e}\dots z_{n}$, we categorize reuse opportunities into three general cases and report their distribution, which is obtained by manually labeling the rejection causes from a user case of summarization. Table 2 summarizes the results, showing that nearly 40% of the cases include reused sequence portions.
  • Missing: the expected sequence includes additional tokens $z_{e}\dots z_{e+\delta}\ (\delta\geq 0)$ not present in the draft, while the remaining draft tokens $x_{e}\dots x_{e+\epsilon}\ (\epsilon\geq 1)$ match the expectation. Thus, the expected sequence is $z_{e}\dots z_{e+\delta}x_{e}\dots x_{e+\epsilon}$. If the combined sequence $z_{e+\delta}x_{e}\dots x_{e+\epsilon}$ does not exist in the database, next retrieval will fail with no drafts generated. This could be recovered by reserving $x_{e}\dots x_{e+\epsilon}$.
  • Synonym: the draft includes a semantically equivalent but lexically different segment $x_{e}\dots x_{e+\gamma-1}\ (\gamma\geq 1)$, replaced in the expected response by $z_{e}\dots z_{e+\delta}$. Similar failure occurs if the post-synonym tokens $x_{e+\gamma}\dots x_{e+\epsilon}$ cannot be retrieved from database using $z_{e+\delta}$. In this case, $x_{e+\gamma}\dots x_{e+\epsilon}$ should be preserved.
  • Redundant: the entire draft $x_{e}\dots x_{n}$ is incorrect and fully diverges from the expected path. This is the only case where discarding the sequence is appropriate.
Reuse methodology. To leverage reusable tokens in the first two cases, CoordGen aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
Enumerating all possible reuse strategies is of $O(N^{3})$ complexity, which is not suitable for online scheduling. Therefore, CoordGen adopts a confidence-based and length-first reuse strategy, assuming that drafts aligned with model predictions are more likely to be accepted in future steps. Therefore, we identify and retain the longest segment of $\mathbf{x}$ such that $x_{i}=y_{i}$ for all $i\in[e+\gamma,e+\epsilon]$, by solving:
(3) $\gamma,\epsilon=\underset{\gamma,\epsilon}{\text{argmax}}(\epsilon-\gamma)\quad\text{s.t. }y_{i}=x_{i},\ \forall\ i\in[e+\gamma,e+\epsilon]$
Each reused segment is reserved for at least one iteration, and is discarded when the cumulative draft length of current decoding step exceeds a predefined threshold.
By maximizing draft length with potentially acceptable tokens, this reuse mechanism directly enhances NPU utilization and reduces verification times. The overhead of CoordGen’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.
4. Evaluation
4.1. Implementation and Setup
Implementation. We implement CoordGen  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code. We reuse the NPU backend of mllm, including neural kernels and memory managers, to perform prefill and draft verification. CoordGen inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities. Our modular design allows easy porting to other mobile LLM inference frameworks that face NPU inefficiencies during decoding. To enable NPU-optimized speculative decoding, we augment mllm with customized kernels and sampling strategies. In order to accelerate drafting, we implement a suffix automaton to identify potential suffix matches (Mohri et al., 2009). To support NPU graph switching, we integrate the graph saving and loading APIs provided by the Qualcomm QNN SDK (Qualcomm, 2025a).
Hardware. We evaluate CoordGen on three representative smartphones with Qualcomm SoCs: Redmi K60 Pro (Snapdragon 8 Gen 2, Android 13, 12 GB RAM), Redmi K70 Pro (Snapdragon 8 Gen 3, Android 14, 16 GB RAM), and OnePlus 13 (Snapdragon 8 Gen Elite, Android 14, 24 GB RAM). All devices are evaluated under real mobile conditions with the CPU frequency governed by the Android OS’s dynamic voltage and frequency scaling controller.
Datasets. We evaluate four common mobile tasks collected from prior work: document summarization (summary), RAG-based question answering (rag), UI automation (ui) and automatic message reply (tweet(Xia et al., 2024; Xing et al., 2024; Salemi et al., 2023).
Models. We use Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct and LLaMA3.2-3B-Instruct (Qwen et al., 2025; Grattafiori et al., 2024), exploring the effect of different model size.
Baselines. We compare CoordGen against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation. CoordGen, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly. EdgeLLM (Xu et al., 2025a) optimizes model-based SD mainly to alleviate memory constraints. With sufficient memory, its performance reduces to that of EAGLE and does not address the memory-bound issue of decoding, thus making it an uninformative baseline for our settings. . All baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (Xu et al., 2025b; Chen et al., 2025a; Li et al., 2024a; Xue et al., 2024). We categorize these baselines by whether they use CPU for decoding or apply speculative decoding:
  • (1)NPU vanilla: use NPU for both prefill and decoding222mllm (Xu et al., 2025b) and other on-device LLM engines  (Chen et al., 2025a; Xue et al., 2024) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality. CoordGen follows this pattern even in the NPU vanilla mode.;
  • (2)CPU vanilla: use NPU for prefill and CPU for decoding, adopted by recent frameworks such as mllm and HeteroLLM (Xu et al., 2025b; Chen et al., 2025a);
  • (3)NPU-SAM: apply retrieval-based speculative decoding (SAM (Hu et al., 2024)) to NPU vanilla. SAM drafts from both historical and local contexts, serving as a standard retrieval-based baseline without mobile-specific optimization;
  • (4)NPU-SAM: similar to (3) but decoding on the CPU;
  • (5)NPU-EAGLE: apply model-based speculative decoding (EAGLE-2 (Li et al., 2024b)) to NPU vanilla. EAGLE-2 is the SOTA model-based SD approach that uses both tokens and hidden states for drafting through a pretrained transformer-like model;
  • (6)CPU-EAGLE: similar to (5) but decoding on the CPU.
Since EAGLE-2 requires a separate draft model that is unavailable for selected LLMs, baselines (5) and (6) are simulated by estimating the expected speed via amortizing the forward costs of the target and draft model. We approximate the draft model as a single transformer layer with a head layer of the target LLM, assuming a fixed maximum acceptance ratio of 5.5, as reported in its paper (Li et al., 2024b).
Metrics and Configuration. We evaluate average per-token latency, energy consumption and peak memory. We exclude accuracy comparison as CoordGen achieves lossless generation (Leviathan et al., 2023a). Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms. The Redmi K70 Pro is excluded from energy results as the device lacks root access. Experiments are repeated three times and we report the average numbers.
4.2. Overall Performance
Per-Token Latency. CoordGenis faster and achieves improvements ranging 1.06–3.81×, as shown in Figure 10.
Figure 10. Per-token latency (ms/token) comparison on different datasets, devices and models. :
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), CoordGen consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×. These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs. Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is memory-bound which is preferred by SD. Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
$\bullet$ Compared to baselines with SD, CoordGen consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE. This arises because SAM ignores the ineffectiveness of computing short drafts on NPUs while EAGLE suffers from the latency of running another parametric model. Improvements of applying SD on CPU vanilla are moderate (0.92–1.13× for SAM, 0.56–1.18× for EAGLE), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
Energy Consumption. CoordGen is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), CoordGen significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×. The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs. On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
$\bullet$ Compared to baselines with SD, CoordGen consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×). Energy saving of SAM is limited due to frequent verification incurred by rejected drafts. EAGLE suffers from additional computation required by drafter, amplifying energy overhead. For instance, for Qwen2.5 0.5B, CoordGen reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
Comparison Across Devices. We find that NPU computation on Redmi K70 Pro is much slower than on other devices, while Redmi K60 Pro and Oneplus 13 demonstrate relatively similar performance patterns. This results in less profound latency improvements on Redmi K70 Pro (e.g., 1.14–2.91× compared to NPU vanilla and 1.13–1.64× compared to CPU vanilla) as compute becomes the bottleneck. In terms of energy consumption, Oneplus 13 consumes more energy than Redmi K60 Pro, e.g., average 53.85 J and 56.52 J per request for Qwen2.5 1.5B.
Figure 11. Energy consumption (J) comparison on different datasets, devices and models. :
4.3. Memory Overhead
Figure 12. Peak memory comparison (GB) on different devices and models. :
As shown in Figure 12, CoordGen maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations. For example, CoordGen requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB. This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
The memory overhead of CoordGen primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory. In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that CoordGen delivers superior efficiency without sacrificing memory scalability.
4.4. Latency Breakdown
To analyze the runtime characteristics and latency overhead of CoordGen, we decompose the end-to-end inference latency into prefill, decoding and potential overhead. Measurements are conducted on Oneplus 13 under the summary benchmark, with results shown in Figure 13.
Figure 13. Breakdown of inference latency (ms). :
The results show that CoordGen significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead. Specifically, decoding of NPU vanilla baseline consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization. CoordGen directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
The prefill overhead is due to sampling tokens from logits for distribution calibration, which costs about 2 ms for each input token. The decoding overhead is caused by draft retrieval and tree construction. Benefited from the low time complexity of suffix automaton, our drafter costs less than 3 ms for each decoding step. Overall, our lightweight techniques incur negligible overhead compared to the overall inference cost.
4.5. Ablation Study
Effectiveness of three techniques. Experiments are conducted on Redmi K60 Pro using the summary benchmark across four baselines: (1) NPU vanilla (NPU), (2) NPU-SAM (+SD), (3) NPU-Graph (+G), and (4) NPU-SAM-Graph (+G+SD). Baselines (1)(2) are identical with those in Section 4.2, while (3)(4) integrate progressive graph scheduling into (1) and (2), respectively. Results are reported in Table 3.
Table 3. Ablation experiments on average per-token latency (ms/token).
Model NPU +G +SD +G+SD Ours
Qwen2.5-1.5B 57.41 47.72 35.04 25.35 15.08
Qwen2.5-0.5B 24.46 23.43 18.75 18.45 7.17
LLaMA3.2-3B 71.71 70.03 47.93 46.25 41.62
We observe that progressive graph scheduling provides consistent speed improvements regardless of whether SD is applied. For example, for Qwen2.5-1.5B, graph scheduling reduces latency from 57.41 ms to 47.72 ms. The results also show that in-context distribution calibration and NPU-optimized draft reuse contribute to significant performance gains across three models, reducing latency from 25.35 ms to 15.08 ms for Qwen2.5-1.5B.
Effectiveness in acceptance ratio. To analyze the effectiveness of techniques relevant to R-SD, we evaluate the acceptance ratio gains applying our techniques to existing methods (PLD (Saxena, 2023), SAM (Hu et al., 2024)). PLD drafts from local contexts only while SAM drafts from both local and historical contexts. Experiments are conducted on a cloud server equipped with A100 GPUs. The results in Table 4 show consistent improvements across benchmarks and models. Our techniques improve the acceptance ratio of PLD by 1.05–1.25×, improve SAM by 1.06–1.24×. These improvements directly translate to higher decoding efficiency, as a higher acceptance ratio reduces the number of tokens that must be recomputed. These results also demonstrate that in-context distribution calibration and draft reuse strategy are orthogonal enhancements with existing R-SD methods, making it effortless to integrate advanced techniques.
Table 4. Ablation on acceptance ratio (average accepted tokens per verification).
Dataset Method Q1.5B Q7B L1B
summary PLD 2.53 1.55 1.42
PLD+Ours 2.91 ($\times$1.15) 1.84 ($\times$1.18) 1.70 ($\times$1.19)
SAM 3.60 1.96 1.68
SAM+Ours 4.21 ($\times$1.17) 2.43 ($\times$1.24) 2.07 ($\times$1.24)
rag PLD 3.19 1.63 1.77
PLD+Ours 3.73 ($\times$1.17) 1.89 ($\times$1.16) 2.11 ($\times$1.19)
SAM 4.42 2.29 2.25
SAM+Ours 5.47 ($\times$1.24) 2.80 ($\times$1.22) 2.66 ($\times$1.19)
ui PLD 2.28 1.64 1.57
PLD+Ours 2.64 ($\times$1.16) 1.73 ($\times$1.05) 1.82 ($\times$1.16)
SAM 3.56 2.44 1.72
SAM+Ours 4.01 ($\times$1.13) 2.59 ($\times$1.06) 2.05 ($\times$1.20)
tweet PLD 2.53 2.06 1.49
PLD+Ours 3.17 ($\times$1.25) 2.52 ($\times$1.22) 1.72 ($\times$1.15)
SAM 3.57 2.98 1.79
SAM+Ours 3.82 ($\times$1.07) 3.70 ($\times$1.24) 2.18 ($\times$1.22)
Q1.5B = Qwen2.5 1.5B, Q7B = Qwen2.5 7B, L1B = LLaMA3.2 1B.
5. Related Work
Mobile LLM inference optimization. The constrained resources of mobile devices require reducing computation to enable efficient on-device LLM inference (Xu et al., 2024; Subramanian et al., 2025; Ma et al., 2024; Thawakar et al., 2024; Xu et al., 2025a). Several approaches exploit different forms of sparsity. Mixture-of-experts leverages activation sparsity by computing only neurons important to the final output (Yi et al., 2025; Chen et al., 2025c; Wu et al., 2025b; Frantar and Alistarh, 2024). Early-exit reduces layer-level computation by skipping non-critical layers (Venkatesha et al., 2025; Fan et al., 2024). Model compression reduces model size by retaining only the most valuable layers (Yu et al., 2024; Smith et al., 2025). Token pruning eliminates redundant tokens to shorten the context (Liu et al., 2025b; Ma et al., 2023). CoordGen is compatible with these techniques, as it requires no architectural modifications to models or inputs.
Mobile NPU offloading. Modern mobile SoCs are increasingly equipped with high-performance NPUs (OPPO, 2024; Samsung, 2024; Huawei, 2024), offering new opportunities for LLM acceleration (Xue et al., 2024; Chen et al., 2025a; Lu et al., 2025; Xu et al., 2025b; Yin et al., 2025; Wei et al., 2025). PowerInfer-V2 (Xue et al., 2024) integrates parameter swapping with NPU offloading to scale inference to larger LLMs. HeteroLLM (Chen et al., 2025a) introduces tensor partitioning strategies tailored to mobile GPUs and NPUs for efficient parallelization. While these studies primarily focus on accelerating prefill, CoordGen complements them by incorporating speculative decoding to optimize both prefill and decoding.
Retrieval-augmented generation. Retrieval-augmented generation (RAG) enhances LLM’s knowledge by retrieving relevant documents from an external database, enabling effective context-aware generation (Lewis et al., 2020). Recent work explores advanced database architectures such as knowledge graph (Edge et al., 2024; Han et al., 2025; Zhang et al., 2025). Other efforts optimize retrieval with adaptive strategies, such as extracting fine-grained information and caching key documents based on user demands (Jeong et al., 2024; Seemakhupt et al., 2024; Liu et al., 2025a). RAG can be employed to augment the context for CAG.
Speculative decoding uses LLM as a verifier to process multiple draft tokens in parallel, shifting memory-bound decoding to compute-bound (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023b). The main difference between these studies is how drafts are generated. Model-based methods employ an auxiliary parametric model, typically a smaller LLM (Li et al., 2024c, b; Cai et al., 2024; Du et al., 2024; Li et al., 2025b). Although these approaches improve the acceptance ratio, they introduce significant drafting overhead (Hu et al., 2024). Retrieval-based methods avoid this cost by using lightweight retrievers for drafting, but often suffer from low acceptance ratios (Saxena, 2023; Somasundaram et al., 2024; Hu et al., 2024; Chen et al., 2025b; Li et al., 2025a; Gritta et al., 2025). CoordGen advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
6. Conclusions
In this paper, we propose CoordGen, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation. CoordGen incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs. Extensive experiments demonstrate CoordGen achieves outstanding and consistent performance gains.
Appendix A NP-hardness Proof of the Graph Scheduling Problem
We prove that the graph scheduling problem described in Section 3.2 is NP-hard via a reduction from the 2-Partition problem (Korf, 1998).
Graph scheduling problem. Given $N$ model blocks that sequentially execute prefill for $C$ chunks, each block $i$ must execute a prefill for every chunk and perform exactly one graph switch from $G^{1}$ to $G^{2}$. Using $G^{2}$ to prefill is slower than using $G^{1}$. Switches can run in parallel with the computation of other blocks, while switches / computation of different blocks cannot overlap with each other. We aim to determine the time to switch each block with an objective to minimize the maximum of the prefill completion time and the switch completion time.
2-Partition problem. Given a set of positive integers $a_{1},\dots,a_{m}$ with sum $S=2B$, determine whether there exists a subset $I\subseteq\{1,\dots,m\}$ such that $\sum_{j\in I}a_{j}=B$.
Construction of the scheduling instance. Given a Partition instance $a_{1},\dots,a_{m}$, we construct a scheduling instance as follows:
  • Let $N=m+1$. Indices $1,\dots,m$ correspond to the $m$ items of the Partition instance, and the block 0 (or equivalently $m+1$) serves as a long block.
  • Let the number of chunks be $C=2$.
  • For each $j=1,\dots,m$ (corresponding to $a_{j}$):
    • Load time: $s_{j}=a_{j}$.
    • Compute time using $G^{2}$: $d_{j}=\varepsilon$, where $\varepsilon>0$ is a very small constant.
    • Compute time using $G^{1}$: $p_{j}<d_{j}$.
  • For the additional block 0:
    • Load time $s_{0}=0$.
    • Compute time using $G^{1}$: $p_{0}=B$.
    • Compute time using $G^{2}$: $d_{0}>p_{0}$.
  • We ask whether there exists a schedule such that the overall latency $T$ is not larger than the prefill completion time.
Key properties. This instance thus satisfy:
  • For each chunk, block 0 follows blocks $1,\dots,m$ and occupies $p_{0}=B$ units of prefill time. Thus, in chunk 1 and chunk 2 we obtain two disjoint windows of length $B$ each. Any switching placed inside one of these windows does not extend the prefill completion time. Conversely, placing a switch outside these windows delays completion beyond $T$.
  • Since all switchings must fit within these two windows of total length $2B$, the only way to complete them before prefill ends is to divide them into two groups of total duration $B$ each. This is exactly the Partition problem.
Correctness.
  • ($\Rightarrow$) If the Partition instance is solvable, i.e., there exists a subset $I$ such that $\sum_{j\in I}a_{j}=B$, then in the scheduling instance we place all switchings for indices in $I$ inside the window of chunk 1 (block 0), and place the rest in chunk 2’s window. Each window contains exactly $B$ switching time, all of which finish before the prefill completion time. Hence, this scheduling satisfies that the overall latency is not larger than the prefill completion time.
  • ($\Leftarrow$) If the scheduling instance admits a feasible schedule with overall latency $\leq T$, then all switchings must be finished before the prefill total time $T$. The only available periods where switchings can be hidden without delaying prefill are the two $B$-length windows (one in each chunk at block 0). Since switchings are serialized, their total time $2B$ must be exactly distributed across the two windows. This requires splitting the switchings into two groups summing to $B$ each, which directly provides a solution to the Partition instance.
    Thus, we have equivalence:
    $\text{Partition is solvable}\;\;\Longleftrightarrow\;\;\text{Scheduling instance is solvable}.$
Conclusion. The reduction can be performed in polynomial time. Since Partition is NP-complete, the prefill scheduling problem is NP-hard.

### v3

@@
-Accelerating Mobile Language Model Generation via Hybrid Context and Hardware Coordination
+Accelerating Mobile Language Model via Speculative Decoding and NPU-Coordinated Execution
 Abstract.
@@
 While recent developments in neural processors have substantially improved the efficiency of prefill on mobile devices, the token-by-token generation process still suffers from high latency and limited hardware utilization due to its inherently memory-bound characteristics.
-This work presents CoordGen, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices.
+This work presents sd.npu, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices.
 The framework introduces three synergistic components: (1) adaptive execution scheduling, which dynamically balances compute graphs between prefill and decoding phases; (2) context-aligned drafting, which improves speculative efficiency through lightweight online calibration to current tasks; and (3) hardware-efficient draft extension, which reuses and expands intermediate sequences to improve processing parallelism and reduce verification cost.
@@
 Motivated by the NPU’s superior energy efficiency and lower contention, this paper raises the question whether NPUs can be effectively utilized to accelerate the decoding stage.
-Figure 1. Workflow of context-augmented generation and comparison between CoordGen and related work :
+Figure 1. Workflow of context-augmented generation and comparison between sd.npu and related work :
 To address this issue, we observe that CAG outputs exhibit strong similarity to both the interaction history and the augmented context (Section 2.2), making them well-suited for retrieval-based speculative decoding (R-SD), a technique originally developed for code generation, and in turn, to NPU acceleration.
@@
 By evaluating multiple drafts in parallel, R-SD shifts decoding from a memory-bound to a compute-bound workload, making it a strong fit for NPU acceleration.
-In this paper, we present CoordGen, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading.
-During prefill, CoordGen employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification.
-Designing CoordGen requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
+In this paper, we present sd.npu, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading.
+During prefill, sd.npu employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification.
+Designing sd.npu requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
 $\bullet$ NPU execution: costly switching overhead for distinct prefill and decoding optimal compute graphs.
@@
 $\bullet$ Progressive Graph Scheduling (Section 3.2).
-CoordGen deploys two specialized graphs and switches between them asynchronously to minimize overhead.
+sd.npu deploys two specialized graphs and switches between them asynchronously to minimize overhead.
 The LLM is partitioned into blocks.
@@
 This design rests on two observations: (i) executing short-input graphs multiple times is equivalent to a single long-input execution, and (ii) graph loading can be overlapped with one-round chunked prefill computation.
-Consequently, CoordGen can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
+Consequently, sd.npu can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
 $\bullet$ In-Context Distribution Calibration (Section 3.3).
-Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, CoordGen uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search.
-In the subsequent retrieval-based drafting stage, CoordGen retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations.
+Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, sd.npu uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search.
+In the subsequent retrieval-based drafting stage, sd.npu retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations.
 Compared with costly precomputation or task-specific fine-tuning, this lightweight procedure incurs negligible overhead.
 $\bullet$ NPU-Optimized Draft Reuse (Section 3.4).
-CoordGen extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy.
+sd.npu extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy.
 This is motivated by the observation that rejected tokens may still capture correct semantics and be accepted in later decoding steps.
-By reusing such tokens, CoordGen increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
+By reusing such tokens, sd.npu increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
 […]
@@
 $\bullet$ We present the first study of inefficiencies in context-aware generation for mobile tasks, identifying key optimization opportunities of using retrieval-based speculative decoding.
-$\bullet$ We design and implement CoordGen, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
-$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that CoordGen delivers consistent and significant performance gains in diverse mobile scenarios.
+$\bullet$ We design and implement sd.npu, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
+$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that sd.npu delivers consistent and significant performance gains in diverse mobile scenarios.
 2. Background
@@
 3.1. Overview
-Figure 7. Overview of CoordGen framework.:
+Figure 7. Overview of sd.npu framework.:
 Design goal.
-CoordGen accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding.
+sd.npu accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding.
 This entails (1) deploying and scheduling two compute graphs optimized for prefill and decoding, (2) aligning task context with LLM distribution, and (3) enriching drafts with sufficient high-value tokens to fully utilize NPU resources.
-Workflow. Figure 7 illustrates the system architecture and workflow of CoordGen.
+Workflow. Figure 7 illustrates the system architecture and workflow of sd.npu.
   • Offline stage.
-CoordGen partitions the LLM into multiple blocks, each containing several transformer layers.
+sd.npu partitions the LLM into multiple blocks, each containing several transformer layers.
 It profiles the load and compute latency of each block through warm-up to determine scheduling strategy.
   • Online stage.
-Upon receiving a request, CoordGen fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph.
+Upon receiving a request, sd.npu fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph.
 During prefill, progressive graph scheduling (Section 3.2) incrementally switches the prefill-optimized graph to the decoding-optimized graph with load overlapped by computation.
 After prefill, output logits are used for in-context distribution calibration (Section 3.3), augmenting the database with calibrated contexts.
-CoordGen then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database.
+sd.npu then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database.
 It iteratively constructs a drafting tree and verifies it via a confidence-based strategy to identify and reuse valuable tokens (Section 3.4).
 After an end-of-sequence token is predicted, the final response is returned and stored in the database manager as a historical context for future requests.
-Key techniques. CoordGen incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
+Key techniques. sd.npu incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
   • Progressive NPU graph scheduling (Section 3.2) enables asynchronous graph switching by overlapping load and computation across chunked prefill rounds.
@@
 […]
-Switching graph progressively. CoordGen gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
+Switching graph progressively. sd.npu gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
   • Graph equivalence by repetition: executing $G^{2}$ multiple times on shorter inputs yields the same result as one execution of $G^{1}$ on longer inputs, allowing early switching.
-As CoordGen maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
+As sd.npu maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
   • Chunk-level scheduling: multiple block computation can be overlapped with a single block loading.
@@
 Figure 8 illustrates progressive scheduling with $N=6$ and two input chunks.
-We suppose that CoordGen decides to overlap loading a block by two blocks of computation based on the profiling results.
+We suppose that sd.npu decides to overlap loading a block by two blocks of computation based on the profiling results.
 Initially, all blocks are loaded with $G^{1}$.
-CoordGen computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$.
+sd.npu computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$.
 Then $\operatorname{P}_{3},\operatorname{P}_{4}$ execute while $G^{2}$ loads for $\operatorname{P}_{5}$.
@@
 Problem Formulation.
-Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, CoordGen aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round.
+Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, sd.npu aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round.
 The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
@@
 This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (Korf, 1998) (proofed in Appendix A).
-CoordGen addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute.
+sd.npu addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute.
 This strategy enables fully masking the load overhead by computation to minimize switching overhead.
-CoordGen finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results.
+sd.npu finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results.
 To hide load latency, the sum of the next $k$ blocks’ compute time must be larger than the load overhead of the last block not switched, formulated as:
 (1) $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})\;\leq\;\sum_{\ell=0}^{k-1}\widehat{T}_{\text{compute}}(\operatorname{P}_{i+\ell};G^{x}_{i+\ell}).$
-CoordGen greedily chooses the minimal feasible $k$ that satisfies Equation 1:
+sd.npu greedily chooses the minimal feasible $k$ that satisfies Equation 1:
 (2) $k^{\star}\;=\;\min\Big\{k\geq 1\;\big|\;\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})\leq\sum_{\ell=0}^{k-1}\widehat{T}_{\text{compute}}(\operatorname{P}_{i+\ell};G^{x}_{i+\ell})\Big\}.$
-Then CoordGen executes $\operatorname{P}_{i},\dots,\operatorname{P}_{i+k^{\star}-1}$ under current graphs while loading $G^{2}$ for $\operatorname{P}_{j}$.
+Then sd.npu executes $\operatorname{P}_{i},\dots,\operatorname{P}_{i+k^{\star}-1}$ under current graphs while loading $G^{2}$ for $\operatorname{P}_{j}$.
 This process repeats until all blocks are switched.
@@
 Workflow.
-After prefill, CoordGen first collects predicted logits for all positions.
-Then CoordGen builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
+After prefill, sd.npu first collects predicted logits for all positions.
+Then sd.npu builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
 […]
@@
 Although these gaps cause rejection due to token-level misalignment, the rejected tokens include correct semantics and may be accepted in the next verification.
-CoordGen identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
+sd.npu identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
 Table 2.
@@
 Reuse methodology.
-To leverage reusable tokens in the first two cases, CoordGen aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
+To leverage reusable tokens in the first two cases, sd.npu aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
 Enumerating all possible reuse strategies is of $O(N^{3})$ complexity, which is not suitable for online scheduling.
-Therefore, CoordGen adopts a confidence-based and length-first reuse strategy, assuming that drafts aligned with model predictions are more likely to be accepted in future steps.
+Therefore, sd.npu adopts a confidence-based and length-first reuse strategy, assuming that drafts aligned with model predictions are more likely to be accepted in future steps.
 Therefore, we identify and retain the longest segment of $\mathbf{x}$ such that $x_{i}=y_{i}$ for all $i\in[e+\gamma,e+\epsilon]$, by solving:
@@
 By maximizing draft length with potentially acceptable tokens, this reuse mechanism directly enhances NPU utilization and reduces verification times.
-The overhead of CoordGen’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.
+The overhead of sd.npu’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.
 4. Evaluation
@@
 Implementation.
-We implement CoordGen  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code.
+We implement sd.npu  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code.
 We reuse the NPU backend of mllm, including neural kernels and memory managers, to perform prefill and draft verification.
-CoordGen inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities.
+sd.npu inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities.
 Our modular design allows easy porting to other mobile LLM inference frameworks that face NPU inefficiencies during decoding.
@@
 Hardware.
-We evaluate CoordGen on three representative smartphones with Qualcomm SoCs: Redmi K60 Pro (Snapdragon 8 Gen 2, Android 13, 12 GB RAM), Redmi K70 Pro (Snapdragon 8 Gen 3, Android 14, 16 GB RAM), and OnePlus 13 (Snapdragon 8 Gen Elite, Android 14, 24 GB RAM).
+We evaluate sd.npu on three representative smartphones with Qualcomm SoCs: Redmi K60 Pro (Snapdragon 8 Gen 2, Android 13, 12 GB RAM), Redmi K70 Pro (Snapdragon 8 Gen 3, Android 14, 16 GB RAM), and OnePlus 13 (Snapdragon 8 Gen Elite, Android 14, 24 GB RAM).
 All devices are evaluated under real mobile conditions with the CPU frequency governed by the Android OS’s dynamic voltage and frequency scaling controller.
@@
 Baselines.
-We compare CoordGen against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation.
-CoordGen, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly.
+We compare sd.npu against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation.
+sd.npu, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly.
 EdgeLLM (Xu et al., 2025a) optimizes model-based SD mainly to alleviate memory constraints.
@@
   • (1)NPU vanilla: use NPU for both prefill and decoding222mllm (Xu et al., 2025b) and other on-device LLM engines  (Chen et al., 2025a; Xue et al., 2024) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality.
-CoordGen follows this pattern even in the NPU vanilla mode.;
+sd.npu follows this pattern even in the NPU vanilla mode.;
   • (2)CPU vanilla: use NPU for prefill and CPU for decoding, adopted by recent frameworks such as mllm and HeteroLLM (Xu et al., 2025b; Chen et al., 2025a);
@@
 We evaluate average per-token latency, energy consumption and peak memory.
-We exclude accuracy comparison as CoordGen achieves lossless generation (Leviathan et al., 2023a).
+We exclude accuracy comparison as sd.npu achieves lossless generation (Leviathan et al., 2023a).
 Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms.
@@
 4.2. Overall Performance
-Per-Token Latency. CoordGenis faster and achieves improvements ranging 1.06–3.81×, as shown in Figure 10.
+Per-Token Latency. sd.npuis faster and achieves improvements ranging 1.06–3.81×, as shown in Figure 10.
 Figure 10. Per-token latency (ms/token) comparison on different datasets, devices and models. :
-$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), CoordGen consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×.
+$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×.
 These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs.
@@
 Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
-$\bullet$ Compared to baselines with SD, CoordGen consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE.
+$\bullet$ Compared to baselines with SD, sd.npu consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE.
 This arises because SAM ignores the ineffectiveness of computing short drafts on NPUs while EAGLE suffers from the latency of running another parametric model.
 Improvements of applying SD on CPU vanilla are moderate (0.92–1.13× for SAM, 0.56–1.18× for EAGLE), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
-Energy Consumption. CoordGen is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
-$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), CoordGen significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×.
+Energy Consumption. sd.npu is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
+$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×.
 The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs.
 On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
-$\bullet$ Compared to baselines with SD, CoordGen consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×).
+$\bullet$ Compared to baselines with SD, sd.npu consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×).
 Energy saving of SAM is limited due to frequent verification incurred by rejected drafts.
 EAGLE suffers from additional computation required by drafter, amplifying energy overhead.
-For instance, for Qwen2.5 0.5B, CoordGen reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
+For instance, for Qwen2.5 0.5B, sd.npu reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
 Comparison Across Devices.
@@
 Figure 12. Peak memory comparison (GB) on different devices and models. :
-As shown in Figure 12, CoordGen maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations.
-For example, CoordGen requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB.
+As shown in Figure 12, sd.npu maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations.
+For example, sd.npu requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB.
 This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
-The memory overhead of CoordGen primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory.
-In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that CoordGen delivers superior efficiency without sacrificing memory scalability.
+The memory overhead of sd.npu primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory.
+In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that sd.npu delivers superior efficiency without sacrificing memory scalability.
 4.4. Latency Breakdown
-To analyze the runtime characteristics and latency overhead of CoordGen, we decompose the end-to-end inference latency into prefill, decoding and potential overhead.
+To analyze the runtime characteristics and latency overhead of sd.npu, we decompose the end-to-end inference latency into prefill, decoding and potential overhead.
 Measurements are conducted on Oneplus 13 under the summary benchmark, with results shown in Figure 13.
 Figure 13. Breakdown of inference latency (ms). :
-The results show that CoordGen significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead.
+The results show that sd.npu significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead.
 Specifically, decoding of NPU vanilla baseline consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization.
-CoordGen directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
+sd.npu directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
 The prefill overhead is due to sampling tokens from logits for distribution calibration, which costs about 2 ms for each input token.
@@
 Token pruning eliminates redundant tokens to shorten the context (Liu et al., 2025b; Ma et al., 2023).
-CoordGen is compatible with these techniques, as it requires no architectural modifications to models or inputs.
+sd.npu is compatible with these techniques, as it requires no architectural modifications to models or inputs.
 Mobile NPU offloading.
@@
 HeteroLLM (Chen et al., 2025a) introduces tensor partitioning strategies tailored to mobile GPUs and NPUs for efficient parallelization.
-While these studies primarily focus on accelerating prefill, CoordGen complements them by incorporating speculative decoding to optimize both prefill and decoding.
+While these studies primarily focus on accelerating prefill, sd.npu complements them by incorporating speculative decoding to optimize both prefill and decoding.
 Retrieval-augmented generation.
@@
 Retrieval-based methods avoid this cost by using lightweight retrievers for drafting, but often suffer from low acceptance ratios (Saxena, 2023; Somasundaram et al., 2024; Hu et al., 2024; Chen et al., 2025b; Li et al., 2025a; Gritta et al., 2025).
-CoordGen advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
+sd.npu advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
 6. Conclusions
-In this paper, we propose CoordGen, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation.
-CoordGen incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs.
-Extensive experiments demonstrate CoordGen achieves outstanding and consistent performance gains.
+In this paper, we propose sd.npu, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation.
+sd.npu incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs.
+Extensive experiments demonstrate sd.npu achieves outstanding and consistent performance gains.
 Appendix A NP-hardness Proof of the Graph Scheduling Problem

### v4

@@
 Abstract.
-Enhancing on-device large language models (LLMs) with contextual information from local data enables personalized and task-aware generation, powering use cases such as intelligent assistants and UI agents.
-While recent developments in neural processors have substantially improved the efficiency of prefill on mobile devices, the token-by-token generation process still suffers from high latency and limited hardware utilization due to its inherently memory-bound characteristics.
-This work presents sd.npu, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices.
-The framework introduces three synergistic components: (1) adaptive execution scheduling, which dynamically balances compute graphs between prefill and decoding phases; (2) context-aligned drafting, which improves speculative efficiency through lightweight online calibration to current tasks; and (3) hardware-efficient draft extension, which reuses and expands intermediate sequences to improve processing parallelism and reduce verification cost.
-Experiments on multiple smartphones and representative workloads show consistent improvements of up to 3.8× in generation speed and 4.7× in energy efficiency compared with existing mobile inference solutions.
-Component-level analysis further validates the contribution of each optimization.
+Performing Retrieval-Augmented Generation (RAG) directly on mobile devices is promising for data privacy and responsiveness but is hindered by the architectural constraints of mobile NPUs.
+Specifically, current hardware struggles with the variable workloads intrinsic to RAG: the transition between processing extensive contexts and generating tokens incurs significant overhead due to static graph constraints, while the memory-bound generation phase leaves computational resources underutilized.
+[…]
 1. Introduction
-Recent advances in large language models (LLMs) (Brown et al., 2020; OpenAI et al., 2024; Touvron et al., 2023; Guo et al., 2024), rising concerns over privacy, and rapid progress in hardware have fueled interest in on-device LLM.
-In particular, the integration of neural processing units (NPUs) (Qualcomm, 2024, 2025b) makes it feasible to run LLMs directly on mobile devices (Google, 2024; Huawei, 2024; Samsung, 2024; Apple, 2024a), enabling applications such as voice assistants and UI automation (Apple, 2024b; AI, 2024; Wen et al., 2024; Lee et al., 2024).
-To extend LLMs’ capabilities in handling complex tasks, context-augmented generation (CAG) has emerged as a key technique (Wang and Chau, 2024; Ouyang et al., 2025).
-As shown in Figure 1, CAG augments a user’s query with task-relevant contexts before inference to strengthen the knowledge grounding of LLM.
-With CAG, LLMs can, for example, summarize documents using user-uploaded files, generate responses by retrieving and integrating information from the chat history, and act as UI agents by reasoning over current UI screenshots captured from the operating system, significantly expanding their functionality.
-Despite these benefits, on-device CAG still has high end-to-end latency, primarily from decoding (accounting for 73.5% of total inference, details in Section 2.1).
-When diving into the reasons, we find that the prefill bottleneck is largely eliminated by NPU-based chunked prefill (Xu et al., 2025b; Chen et al., 2025a; Xue et al., 2024) (only 2 s for 1–2K tokens context).
-In contrast, the decoding stage remains memory-bound, a scenario poorly suited for NPUs.
-Therefore, most existing work resorts to CPU execution, leading to high latency.
-Motivated by the NPU’s superior energy efficiency and lower contention, this paper raises the question whether NPUs can be effectively utilized to accelerate the decoding stage.
-Figure 1. Workflow of context-augmented generation and comparison between sd.npu and related work :
-To address this issue, we observe that CAG outputs exhibit strong similarity to both the interaction history and the augmented context (Section 2.2), making them well-suited for retrieval-based speculative decoding (R-SD), a technique originally developed for code generation, and in turn, to NPU acceleration.
-In many CAG tasks, generated content is semantically close to previously produced tokens and to the augmented context: for example, summarization and automated email reply often reuse spans from user-uploaded documents or earlier emails, while UI agents repeatedly emit fixed commands (e.g., “click set clock button”).
-These characteristics align naturally with R-SD, which proceeds in two stages: (1) drafting, where a retriever proposes candidate token sequences by selecting similar fragments from the augmented context or a local store (e.g., a code repository); and (2) verifying, where the LLM compares these drafts against its own next-token predictions and corrects discrepancies (see Section 2.2).
-By evaluating multiple drafts in parallel, R-SD shifts decoding from a memory-bound to a compute-bound workload, making it a strong fit for NPU acceleration.
-In this paper, we present sd.npu, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading.
-During prefill, sd.npu employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification.
-Designing sd.npu requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
-$\bullet$ NPU execution: costly switching overhead for distinct prefill and decoding optimal compute graphs.
-Prefill favors large-shape graphs with long sequence inputs which degrade decoding performance, while decoding benefits from small-shape graphs that underperform for prefill.
-Since the NPU only supports inference on static shapes, achieving optimal end-to-end performance thus requires switching between prefill-optimal and decoding-optimal compute graphs, incurring a significant performance overhead.
-For instance, our preliminary evaluation on Qwen2.5-1.5B model shows this switching overhead incurs a 94.9% delay of end-to-end latency.
-$\bullet$ Draft construction: task context and LLM outputs follow different lexical distributions, yielding a low acceptance ratio.
-In R-SD, the LLM only accepts tokens that exactly match its own prediction.
-Our preliminary evaluation shows that drafts constructed from the task context often have high semantic similarity to the target output yet exhibit notable lexical gaps (e.g., normalized Levenshtein distance >0.11>0.11 for Qwen2.5-1.5B).
-As a result, LLM often accepts only few tokens per draft, falling back to token-by-token decoding and underscoring the need for lexical distribution calibration.
-$\bullet$ Draft verification: short or absent drafts fail to saturate NPU computation.
-Mobile NPUs excel at processing large batches or long sequences due to their weight-stationary design where short inputs obtain fewer benefits (Jouppi et al., 2017; Chen et al., 2025a).
-However, due to strict suffix-matching, existing R-SD methods often generate only a few candidate tokens or none at all.
-Such short sequences cannot fully exploit the NPU’s weight-stationary design, resulting in low verification throughput.
-For example, our experiments show that over 75% of the drafts fall below length 8, which is far below the verification-optimal length (64) and leads to 50% lower throughput.
-To address these challenges, we propose three techniques:
-$\bullet$ Progressive Graph Scheduling (Section 3.2).
-sd.npu deploys two specialized graphs and switches between them asynchronously to minimize overhead.
-The LLM is partitioned into blocks.
-During each round of chunked prefill, we switch to a decoding-optimized block graph while prefill computation continues in parallel.
-The switched decoding graph is invoked multiple times within the prefill stage to ensure result correctness.
-This design rests on two observations: (i) executing short-input graphs multiple times is equivalent to a single long-input execution, and (ii) graph loading can be overlapped with one-round chunked prefill computation.
-Consequently, sd.npu can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
-$\bullet$ In-Context Distribution Calibration (Section 3.3).
-Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, sd.npu uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search.
-In the subsequent retrieval-based drafting stage, sd.npu retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations.
-Compared with costly precomputation or task-specific fine-tuning, this lightweight procedure incurs negligible overhead.
-$\bullet$ NPU-Optimized Draft Reuse (Section 3.4).
-sd.npu extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy.
-This is motivated by the observation that rejected tokens may still capture correct semantics and be accepted in later decoding steps.
-By reusing such tokens, sd.npu increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
+Retrieval-augmented generation (RAG) has become the de facto standard for grounding Large Language Model (LLM) outputs in user-specific data, enabling a new class of edge applications ranging from personalized document assistants to automated UI agents (usecase:ui_automation; work:mobile_rag1; work:mobile_rag2; work:mobile_rag3).
+As these applications increasingly handle sensitive personal data, RAG workloads are shifting towards on-device execution to ensure privacy and reduce cloud reliance (trend:slm1; trend:slm2; privacy1; privacy2; privacy3).
+In this setting, the retrieval phase is typically fast (e.g., <10<10 ms for local corpora (work:edge_rag; work:webanns)), leaving the computational heavy lifting to the LLM generation phase.
+To sustain always-on RAG capabilities without draining the battery or stalling the user interface, the Neural Processing Unit (NPU) has emerged as the critical hardware accelerator.
+Unlike CPUs and GPUs that are burdened with OS services, application logic, and graphics rendering, NPUs are domain-specific architectures dedicated to tensor operations.
+They offer two distinct advantages: (i) superior energy efficiency, delivering significantly lower energy-per-token for Transformer workloads compared to general-purpose processors (work:mllm; work:heterollm; work:tman); and (ii) performance isolation, allowing background RAG inference to run concurrently with foreground tasks without contending for CPU/GPU resources (work:npu_sparse_attn).
+Driven by these benefits, recent systems have moved towards NPU-only inference (work:tman; work:ttc; work:npu_sparse_attn), executing the entire pipeline on the NPU to eliminate the high synchronization and data movement costs inherent in heterogeneous CPU-NPU co-scheduling (work:mllm; work:heterollm).
+Despite these advantages, NPU-only on-device RAG inference still suffers from long decoding latency, driven by two critical challenges:
+$\bullet$ Static graph architecture vs.
+dynamic RAG phases.
+Mobile NPUs rely on static compute graphs with fixed tensor shapes, buffer layouts, and execution plans.
+Yet RAG inference alternates between two computational extremes: prefill, which processes long, context-heavy prompts (suitable graph for 128/256 tokens input), and decoding, which autoregressively generates tokens one-by-one (suitable graph for 1 token input).
+The gap between these two regimes is too large that no single static graph can efficiently serve both.
+To fit within tight on-device memory budgets, existing systems keep only one graph resident at a time, necessitating an expensive teardown-and-load process when switching phases: unloading the large prefill graph and loading the decoding graph.
+We find that naive graph switching accounts for up to 29% of the decoding latency in on-device RAG settings.
+$\bullet$ Matrix-oriented hardware vs.
+memory-bound decoding.
+While prefill saturates the NPU’s matrix units, decoding is memory-bound, leaving compute resources severely underutilized.
+Speculative Decoding (SD) (work:sd_is_lossless) is a standard remedy that enables parallel decoding through a draft-then-verify schema, but existing model-based SD methods (e.g., EAGLE (work:eagle)) require auxiliary models, leading to high mobile memory consumption and slow speculative inference on NPUs.
+For RAG specifically, retrieval-based SD (work:sam) is theoretically ideal: it drafts candidate tokens directly from context by first-token retrieval matching without requiring the extra cost of auxiliary draft models.
+However, we observe that off-the-shelf retrieval-based SD fails to saturate mobile NPUs due to two inefficiencies.
+To understand why, we introduce the concept of effective NPU utilization and decompose the NPU cycles spent during verification into three categories: (i) Valid: cycles spent verifying tokens that are eventually accepted; (ii) Wasted: cycles spent verifying tokens that are rejected; and (iii) Underutilized: idle capacity in the matrix engines when the workload is too small (low batch size/sequence length).
+Our analysis (§ 3) shows that existing approaches are dominated by the following inefficiencies:
+  1. (i)
+    Lexical Divergence causes Wasted Compute: retrieved context often matches the semantic intent but differs lexically from the LLM’s training distribution.
+This causes the verifier to reject most draft tokens.
+In our measurements, 22% of NPU cycles are wasted verifying drafts that are ultimately rejected.
+  2. (ii)
+    Short Drafts cause Underutilization: NPU matrix units (e.g., Qualcomm HMX (tool:qnn_sdk), Huawei Cube (tool:ascend)) rely on large operational intensity (large batch or sequence length) to hide memory latency and satisfy the minimal tile size (e.g., 32 tokens for HMX).
+However, retrieval-based drafts are often fragmented and short.
+We find that over 75% of drafts are shorter than 8 tokens, whereas mobile NPUs typically require sequences of 32+ tokens to approach peak utilization.
+This mismatch leaves more than 70% of the NPU’s compute capacity underutilized during verification steps.
+Design We present sd.npu, the first system designed to enable efficient, end-to-end NPU offloading for on-device RAG.
+It bridges the gap between the NPU’s requirement for static, massive parallelism and the dynamic-shape, fragmented nature of RAG workloads by incorporating several techniques:
+$\bullet$ Progressive graph scheduling (§ 4.2) addresses the monolithic loading stall caused by deferring graph switching to the final prefill stage.
+Instead of a sudden, blocking switch, we amortize the heavy I/O cost across the entire prefill phase by leveraging two key insights.
+(i) Decomposition: the monolithic LLM is partitioned into fine-grained blocks (subsets of Transformer layers), making the I/O payload of a single block small enough to be masked by several chunked prefill computation.
+(ii) Equivalence: we observe that repeated, iterative execution of a decoding-friendly block graph is mathematically equivalent to a single execution of a prefill-optimized graph.
+Based on these, sd.npu orchestrates a multi-iteration progressive transition.
+Specifically, by partitioning the LLM, we progressively orchestrates the block-by-block switching of the prefill graphs to their decoding-friendly counterparts; each block switching is meticulously designed to be perfectly overlapped with one chunked prefill operation.
+This ensures a seamless and zero-cost transition to the decoding phase.
+After the decoding graph for each block has been loaded, it is executed iteratively to process the corresponding chunk of the prefill input, thereby ensuring functional equivalence and maintaining the correct overall prefill output.
+$\bullet$ NPU-tailored speculative decoding (§ 4.3, 4.4).
+Building upon retrieval-based speculative decoding, sd.npu transforms the ”wasted” and ”underutilized” NPU cycles into ”valid” generation throughput, incorporating two distinct techniques:
+  1. (i)
+    In-context distribution calibration (§ 4.3).
+To minimize wasted compute caused by lexical divergence, we must align the retrieved drafts with the model’s output distribution.
+Cloud-based solutions might perform offline fine-tuning on the knowledge base, but this is infeasible on edge devices due to battery, storage, and privacy constraints.
+Instead, sd.npu adopts a zero-cost, online calibration.
+We observe that during the prefill, the model inherently computes the logits for the provided context.
+sd.npu reuses these existing logits to construct a model-calibrated token tree.
+By drafting from this calibrated tree rather than the raw text, we reduce draft rejection without any additional computation or model training.
+  2. (ii)
+    NPU-optimized draft extension (§ 4.4).
+Minimizing underutilized compute caused by short drafts requires longer draft sequences that can saturate the NPU matrix units.
+Existing SD methods focus solely on the input side (drafting) and discard rejected tokens.
+However, we observe that rejection is often temporary: around 38.5% of tokens rejected in one step become valid in subsequent steps as the context evolves.
+Based on this insight, sd.npu implements a verification-aware draft extension.
+It recycles high-confidence rejected tokens from previous steps to extend the current draft.
+This not only recovers useful information but also artificially lengthens the verification batch, pushing the NPU into a higher-utilization regime.
 […]
-Our contributions are summarized as follows:
-$\bullet$ We present the first study of inefficiencies in context-aware generation for mobile tasks, identifying key optimization opportunities of using retrieval-based speculative decoding.
-$\bullet$ We design and implement sd.npu, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
-$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that sd.npu delivers consistent and significant performance gains in diverse mobile scenarios.
+$\bullet$ Identify and characterize the key inefficiencies of end-to-end NPU offloading for RAG workloads, revealing optimization opportunities via speculative decoding.
+$\bullet$ Design and implement sd.npu, a system that co-optimizes speculative decoding and NPU execution through two novel techniques, bridging the gap between compute-bound LLM decoding and NPU offloading.
+$\bullet$ Conduct comprehensive evaluations across multiple devices, models, and tasks, demonstrating consistent latency and energy efficiency improvements in diverse edge environments.
 2. Background
-2.1. On-device Context-Augmented Generation for Mobile Task
-Context-augmented generation.
-With the rapid development of LLMs and increasing demands for privacy protection, there is a growing trend of deploying large models on mobile devices (Belcak et al., 2025; Navardi et al., 2025; Union, 2021; Kshetri, 2023; Eleftherakis et al., 2024).
-An emerging paradigm, context-augmented generation, leverages both the long-context capability of LLMs and the abundance of private on-device data, showing strong potential for supporting complex tasks such as personal assistant and UI automation (Wang and Chau, 2024; Ouyang et al., 2025; Park et al., 2025).
-Typically, context-augmented generation first exploits additional information, either directly provided by the user or retrieved from relevant sources, to enrich the prompt with task-specific contexts.
-Then, the normal LLM inference procedure is executed, including prefill and decoding stages.
-The enriched prompt can alleviate the limitations of the model’s built-in knowledge and enable it to handle a wider range of concrete tasks.
-For instance, in a summarization task, the user may supply a meeting transcript as context, and the LLM directly generates a concise summary from it.
-In another case, a personalized assistant retrieves a user’s past chat records to draft a reply in the user’s typical style, where the retrieved records provide the key context for generation.
-LLMs on NPUs.
-As modern mobile SoCs increasingly integrate high-performance NPUs (Table 1), recent mobile LLM frameworks aim to offload LLM inference onto NPUs to improve efficiency.
-In practice, NPUs are mainly used for the prefill stage, which benefits from their high throughput on long input sequences, while decoding is still executed on CPUs or GPUs (Xu et al., 2025b; Chen et al., 2025a; Xue et al., 2024).
-Our experiments reveal that mobile CAG still suffers from high latency under this design, with decoding dominating end-to-end performance.
-For example, for Qwen2.5-1.5B (Qwen et al., 2025) running on a OnePlus 13, decoding accounts for 90.2% of latency in question answering (QA) tasks and 87.2% in summarization, even when prefill runs at over 1000 tokens/s (Xu et al., 2025b) (Figure 2).
-Simply offloading decoding to the NPU provides no advantage over CPU execution and often increases overall latency, as decoding is inherently memory-bound, processing one token at a time and leaving the NPU underutilized.
-Figure 2. Comparison of prefill and decoding latency for several CAG tasks. Task details are provided in Section 4.1.
-Table 1. Specifications of representative mobile NPUs.
+2.1. RAG for Mobile Applications
+Retrieval‑augmented generation (RAG) has become an essential technique for on‑device intelligent applications (work:rag; work:mobile_rag1; work:mobile_rag2), such as mobile UI agents and personal assistants.
+By retrieving relevant external documents (e.g., UI information for automation and user archives for personalization) and injecting it into the prompt, RAG improves the accuracy and reliability of LLM‑driven tasks.
+Compared to cloud‑based inference, on‑device RAG offers strong privacy guarantees and consistent responsiveness, making it well suited for scenarios involving sensitive data or real‑time interactions (work:adaptive_rag).
+Recent efforts on retrieval acceleration have largely addressed its overhead, achieving latency of several microseconds for querying on GB-size corpus (work:webanns; work:edge_rag).
+Thus, our focus is the compute burden introduced by LLM inference, as the augmented context significantly lengthens prompts, amplifying the cost of both prefill and decoding (work:rag_amplify_cost1; work:rag_amplify_cost2).
+Fortunately, modern mobile SoCs are increasingly integrating neural processor units (NPUs) with high compute capacity, offering opportunities for hardware acceleration.
+2.2. Mobile NPUs
+Figure 1. RAG on NPU and NPU’s architecture.
+Modern mobile NPUs are specialized accelerators designed to deliver high throughput and energy‑efficient execution for deep neural networks.
+Their compute cores are typically built around matrix computation units, which usually are implemented as systolic arrays with weight‑stationary designs (device:8gen3; device:8gen4).
+As illustrated in Figure 1, systolic arrays preload weight tiles into their processing elements and keep them stationary during computation.
+This weight‑stall paradigm amortizes the cost of weight movement and maximizes reuse across input activations, achieving high efficiency for large matrix multiplications.
+Meanwhile, the fixed systolic array constraints the minimal input dimensions to effectively saturate the processing units, as each tile in the array (e.g., 32×32 FP16 or INT8) corresponds to the fundamental execution unit of the NPU.
+To address this, the NPU runtime compiles operators into sequences of tile‑level matmul instructions, pads and partitions input tensors into tiles such that activations could be streamed through the systolic array with the optimal efficiency.
+Therefore, most mobile NPU SDKs require neural networks to be converted into static compute graphs with fixed tensor shapes, memory bindings, and operator layouts.
+Graph compilation can take several seconds and is generally performed offline (work:mllm).
+At runtime, the NPU executes these static graphs with minimal control‑flow support, which simplifies scheduling overhead by restricting dynamic shape changes and conditional execution.
+2.3. Accelerating RAG on Mobile NPUs
 Vendor SoC NPU Perf. (Tops)
@@
 Perf. = INT8 Performance in Tops.
-These findings are closely tied to the architectural characteristics of mobile NPUs.
-First, their systolic-array with weight-stationary design excels at long input sequences: weights are preloaded once and reused across tokens, greatly reducing memory traffic.
-This property aligns naturally with the prefill stage, which requires lots of computation over long prompts.
-Second, NPUs rely on static compute graphs with precompiled execution plans, where operator arrangements and tensor shapes are fixed in advance for efficiency.
-Naively reusing prefill-optimized graphs for decoding leads to severe inefficiencies.
-As a result, while NPUs eliminate prefill bottlenecks, they bring little benefit to decoding, which thus remains the dominant source of latency and energy consumption in on-device CAG.
-2.2. Opportunity: Retrieval-based Speculative Decoding
-A key observation is that CAG naturally inherits similarity from the request or the augmented context, making it well-suited for retrieval-based speculative decoding (Saxena, 2023; He et al., 2024; Hu et al., 2024; Somasundaram et al., 2024), an acceleration technique exploiting similarity between contexts and outputs.
-We find that mobile CAG tasks exhibit the same property, offering strong opportunities for acceleration via R-SD.
-$\bullet$ History similarity in similar tasks: outputs from similar tasks resemble previous generations because their contexts are similar.
-For example, a UI automation agent often repeats commands like clicking the same button.
-Our evaluation on four UI automation task groups (Wen et al., 2024) with Qwen2.5-1.5B (Qwen et al., 2025) confirms this (Figure 3(a)): predictions within the same task show the highest semantic similarity (0.72), those within the same group moderate similarity (0.61), and across groups low similarity (0.38).
-Cosine similarity is computed using embeddings from all-MiniLM-L6-v2 (Wang et al., 2020).
-$\bullet$ Context similarity at semantic level: Outputs align with the enriched context, which often includes direct clues for response (Wang and Chau, 2024; Ouyang et al., 2025; Park et al., 2025).
-For instance, assistants answering QA queries frequently extract information from documents.
-On a summarization dataset (Xia et al., 2024) with Qwen2.5-1.5B, we compute semantic similarity between context and predictions (via cosine similarity) and lexical overlap (via Levenshtein distance).
-Figure 3(b) shows that 90% of samples exhibit high semantic similarity (>0.76>0.76) but low lexical overlap (>0.11>0.11), reflecting a stylistic gap between context and outputs.
-(a) History
-(b) Context
-Figure 3. Demonstration of similarities in context-augmented generation.
-Figure 4. Workflow of retrieval-based speculative decoding.
-R-SD exploits these similarities through two stages (Figure 4): (1) retrieval-based drafting, where relevant fragments are retrieved from history or context to draft candidate sequences.
-For example, tokens “need to” may retrieve drafts such as “click on” and “find the”; (2) verifying, where the LLM validates these candidates in parallel, accepting matches and discarding mismatches to guarantee correctness (Leviathan et al., 2023a).
-In the example where model expects “find the”, mismatched drafts like “click on” are discarded while the matched draft is accepted, extending the output to “find the Google”.
-The similarities in mobile CAG tasks offer promising draft candidates, enabling R-SD to efficiently generate multiple tokens in parallel.
-R-SD thus substantially raises the computation-to-memory ratio, shifting decoding from memory-bound to compute-bound where NPU excels.
-Parallel to retrieval-based methods, model-based speculative decoding (M-SD) (e.g., EAGLE (Li et al., 2024c, b), Medusa (Cai et al., 2024)) trains auxiliary draft models to generate candidates, incurring extra training, storage, and inference overheads.
-R-SD avoids retraining by directly using contextual information of high similarity, achieving draft latencies under 10ms (Hu et al., 2024).
-Therefore, R-SD is more suitable for resource-constrained devices while M-SD suffers from higher latency, as shown in our evaluations (Section 4.2).
-2.3. Challenges: Aligning Retrieval-based Speculative Decoding with NPU
-Although R-SD offers a promising approach to accelerate CAG with NPU offloading, integrating it with NPU offloading poses several challenges.
+Table 1. Specifications of representative mobile NPUs.
+As mobile SoCs increasingly integrate NPUs (Table 1), leveraging these accelerators for RAG has become a critical optimization direction.
+When deploying RAG systems on edge devices, existing inference engines typically adopt one of two execution strategies:
+$\bullet$ NPU-CPU/GPU co-scheduling (e.g., mllm (work:mllm), HeteroLLM (work:heterollm), PowerInfer‑V2 (work:powerinfer2)): offloading only compute-intensive operations in prefill to NPU, while executing precision-intensive kernels and decoding on CPUs or GPUs.
+$\bullet$ NPU-only execution (e.g., T-MAN (work:tman), test-time-compute NPU (work:ttc) and ExecuTorch (work:executorch)): offloading all computation to NPU through carefully designed quantization strategies.
+Figure 2. Comparison on consumed energy per request across different execution backends (work:mllm; work:heterollm).
+Among these, the NPU-only approach is particularly promising for sustainable RAG acceleration.
+As shown in Figure 2, NPU-only execution reduces energy consumption by eliminating frequent cross‑processor scheduling and by consolidating execution onto the most energy‑efficient compute unit (work:mllm; work:heterollm).
+Furthermore, it also minimizes resource contention with other general-purpose mobile workloads processed by CPU and GPU, such as user interaction handling and rendering (work:npu_sparse_attn).
+Therefore, this paper focuses on addressing the challenges within the NPU-only acceleration path.
+However, realizing the full potential of NPU-only execution remains challenging.
+Prior work (e.g., test-time-compute NPU (work:ttc), T-MAN (work:tman)) primarily relies on aggressive quantization strategies (work:decdec), which necessitate invasive changes to NPU kernels and proprietary SDKs.
+Moving beyond kernel-centric approaches, we aim to optimize RAG performance through system-level innovations.
+Our optimizations address the scheduling and memory inefficiencies inherent in NPU runtimes and are designed to be orthogonal to specific quantization schemes, enabling additive performance gains atop these kernel-level baselines.
+3. Challenges of RAG on NPU
+Figure 1 depicts the NPU-based RAG workflow.
+Excluding the negligible retrieval overhead, execution comprises two distinct phases: compute-bound prefill and memory-bound decoding.
+During prefill, the variable-length context is partitioned into fixed-size chunks (e.g., 256 tokens) to accommodate the NPU’s static shape requirement (work:mllm; work:heterollm).
+However, this execution paradigm exposes two critical inefficiencies on resource-constrained mobile devices: (i) graph switching vs.
+padding dilemma (§ 3.1): due to strict memory constraints, mobile runtimes can typically reside only one active static graph.
+Given the drastic shape disparity between prefill (e.g., 256 tokens) and decoding (1 token), the system faces a prohibitive trade-off: either switch graphs, which incurs high latency from model reloading, or reuse the prefill graph via padding, which results in massive computation waste.
+(ii) NPU underutilization in decoding (§ 3.2): the decoding phase fails to saturate the NPU’s massive matrix parallelism.
+A typical widely-used technique to alleviate this by increasing concurrency is speculative decoding (SD) (work:sd_is_lossless; work:eagle2; work:rest).
+However, we find traditional SD still fails to saturate the NPU’s systolic array architecture due to a mismatch with the hardware cost model.
+Crucially, leaving NPU compute resources significantly idle compared to the dense prefill phase.
+3.1. Overhead of Static Graph Management
+[…]
 Figure 3. Trade-off between prefill and decoding with fixed-size compute graphs, highlighting the importance to deploy specialized graphs for each stage.
-$\bullet$ NPU execution: static compute graph design hinders performance improvement for both prefill and decoding.
-Static compute graph constraints fixed input and output tensor shapes during execution.
-To support variable-length sequences in LLM inference, existing systems adopt a chunking strategy, i.e., splitting inputs into fixed-size segments (Xu et al., 2025b; Chen et al., 2025a).
-However, a unified chunk size for both prefill and decoding yields suboptimal performance.
-Figure 5 compares prefill and speculative decoding speed using different compute graphs on a Redmi K70 Pro smartphone, with a fixed acceptance ratio of 4.2 (based on evaluation results, see Section 4.5).
-Results show that compute graphs with larger chunk size improve prefill throughput but hurt decoding performance.
-For instance, despite a fast prefill speed (¡8 ms/token) of Qwen2.5 1.5B using graph with chunk size 32, the decoding speed using the same graph drops to about 160 ms/token.
-This tradeoff necessitates specialized compute graphs to improve performance for both prefill and decoding.
-However, on resource-constrained devices, naively maintaining multiple graphs in memory is costly, highlighting the needs of lightweight graph scheduling algorithms.
-Figure 6. Comparison of verification speed with different draft lengths, and the length distribution of drafts generated by existing R-SD method (Saxena, 2023) (histogram).
-$\bullet$ Draft construction: divergence between task context and LLM output reduces the efficiency of speculative decoding.
-As discussed in Section 2.1, contexts and predictions often exhibit lexical gaps (i.e., normalized Levenshtein distance >0.11>0.11).
-These gaps result in frequent rejection of retrieved drafts, despite their success in capturing the similar semantics as LLM.
-This degrades decoding to a token-by-token process where NPUs offer little speedup due to underutilization.
-Lexical gaps arise from divergence between the task context and the LLM’s distribution, necessitating alignment of the two distributions.
-$\bullet$ Draft verification: existing R-SD methods underutilize NPUs’ weight-stationary design due to drafting short sequences.
-The design of mobile NPUs, while ideal for compute-intensive operations such as prefill, is less effective when input sequences become short in the decoding stage.
-Existing R-SD methods perform retrieval by strict suffix matching between context and output (Somasundaram et al., 2024; He et al., 2024; Hu et al., 2024), which frequently results in short or empty drafts.
-Figure 6 demonstrates the influence of draft length on verification speed and the length distribution of generated draft on a summarization dataset (Xia et al., 2024) .
-Results show that over 90% of drafts fall below the NPU-unfriendly length, significantly reducing the verification speed as NPU computing capacity is underutilized.
-This results in a similar speed to CPU (shown in the horizontal line).
-For instance, the verification speed of Qwen2.5 1.5B drops from 5.7 ms/token to 51.1 ms/token when the draft length decreases from 32 to 1.
+Figure 4. Latency breakdown of RAG decoding phase.
+Consequently, runtimes are compelled to perform graph switching, i.e., unloading the prefill graph and reloading a decoding-specialized graph.
+However, this operation is far from free.
+Our breakdown of decoding latency (Figure 4) reveals that graph switching alone accounts for 29.0% of the total time for RAG tasks with 1K context.
+Naive switching blocks execution and saturates memory bandwidth, while keeping both graphs resident is infeasible due to the tight RAM budget of mobile devices.
+In contrast, CPU decoding avoids this overhead but suffers from significantly lower raw compute capability.
+[…]
+3.2. Ineffectiveness of Speculative Decoding on Mobile NPUs
+Figure 5. Comparison of verification speed with different draft lengths, and the length distribution of drafts generated by existing SD method (work:pld) (histogram).
+Figure 6. Workflow of speculative decoding and an example of existing inefficiencies: a large proportion of computation is wasted or underutilized.
+Besides graph scheduling challenge, the NPU‑only decoding phase faces another bottleneck: memory bandwidth and NPU underutilization.
+Generating a single token on the NPU is memory-bound due to low arithmetic intensity.
+In this regime, the NPU’s massive matrix array is starved, degrading throughput to CPU levels while consuming the fixed overhead of NPU kernel invocation.
+To amortize these overheads and saturate the NPU, the workload must be transformed from sequential vector operations into parallel matrix operations.
+Speculative decoding (work:sd_is_lossless) is a widely adopted method to effect this transformation.
+SD uses a lightweight model (i.e., model-based methods (work:medusa; work:eagle)) or database (i.e., retrieval-based methods (work:pld; work:rest)) to generate multiple draft tokens and verifies them in parallel.
+In verification, only correct tokens are accepted and any tokens after the first mismatch is discarded (Figure 6a).
+This draft-then-verify schema allows SD to convert memory-bound decoding into compute-bound verification while maintaining lossless generation quality (work:sd_is_lossless).
+Among different approaches, model-based SD is infeasible for mobile devices as it requires a separate draft model (often >1>1GB), which competes for scarce RAM and compute capacity with the main LLM and the OS.
+Instead, for RAG workloads, the retrieved documents intrinsically contain the information needed for generation.
+Using the context as a draft source (work:sam) incurs negligible overhead, making it the ideal candidate for mobile RAG.
+Consequently, this paper adopts retrieval-based speculative decoding as the foundational mechanism for mobile acceleration.
+Breakdown of SD inefficiency.
+Despite its theoretical suitability, we find that off-the-shelf algorithms fail to accelerate NPU inference.
+To understand this inefficiency, we decompose NPU computation during verification into three parts: (i) valid, resources spent on accepted drafts; (ii) wasted, resources spent on rejected drafts; and (iii) underutilized, the NPU’s idle capacity when the workload is too small.
+As show in Figure 6b, our measurements reveal that wasted (22%) and underutilized portions (75%) dominate the use of compute resources, leaving the effective throughput surprisingly low.
+Figure 7. Effective throughput comparison. For NPU, throughput is dictated by accept length.
+Root cause: constant-cost trap.
+This inefficiency stems from the unique performance characteristics of mobile NPUs.
+Unlike GPUs where cost scales linearly, NPU verification latency remains nearly constant up to a hardware saturation point $M$ (Figure 7).
+In this regime, decoding throughput is dictated solely by the total accepted length per step.
+To achieve speedup, the system must maximize the number of accepted tokens within this constant-time window.
+However, existing retrieval-based methods fail to exploit this characteristic.
+For instance, as shown in § 5.5, PLD (work:pld) achieves an average accepted length of only 2.53 tokens per step, resulting in less speedup.
+This failure manifests in two specific ways:
+Challenge#2.1: wasted computation due to low draft acceptance.
+To preserve lossless generation, SD rejects drafts that lexically mismatch the base model’s predictions, even if they are semantically aligned.
+In RAG workloads, this lexical divergence is common, causing frequent rejections.
+Since the NPU incurs a near-constant high cost for each verification step regardless of the outcome, every rejection squanders a full pass of NPU computation.
+This directly contributes to a high percentage of wasted computation.
+Challenge#2.2: underutilized computation due to short workloads.
+The NPU’s constant-cost model means that peak efficiency is only achieved with long drafts that saturate its compute units.
+However, existing SD methods for RAG often produce very short drafts.
+Empirically, over 90% of drafts contain fewer than 32 tokens (Figure 5).
+Verifying such short sequences leaves most of the NPU’s matrix units idle, paying the high cost of a full verification step for only a few tokens.
+This severely throttles effective throughput, sometimes making NPU even slower than CPU.
+Together, these observations motivate the need for an new SD algorithm explicitly co-designed with RAG workloads and NPU execution characteristics.
 4. Methodology
 4.1. Overview
-Figure 7. Overview of sd.npu framework.:
+Figure 8. Overview of sd.npu.
 Design goal.
-sd.npu accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding.
-This entails (1) deploying and scheduling two compute graphs optimized for prefill and decoding, (2) aligning task context with LLM distribution, and (3) enriching drafts with sufficient high-value tokens to fully utilize NPU resources.
-Workflow. Figure 7 illustrates the system architecture and workflow of sd.npu.
-  • Offline stage.
-sd.npu partitions the LLM into multiple blocks, each containing several transformer layers.
-It profiles the load and compute latency of each block through warm-up to determine scheduling strategy.
-  • Online stage.
-Upon receiving a request, sd.npu fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph.
-During prefill, progressive graph scheduling (Section 3.2) incrementally switches the prefill-optimized graph to the decoding-optimized graph with load overlapped by computation.
-After prefill, output logits are used for in-context distribution calibration (Section 3.3), augmenting the database with calibrated contexts.
-sd.npu then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database.
-It iteratively constructs a drafting tree and verifies it via a confidence-based strategy to identify and reuse valuable tokens (Section 3.4).
-After an end-of-sequence token is predicted, the final response is returned and stored in the database manager as a historical context for future requests.
-Key techniques. sd.npu incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
-  • Progressive NPU graph scheduling (Section 3.2) enables asynchronous graph switching by overlapping load and computation across chunked prefill rounds.
-  • In-context distribution calibration (Section 3.3) aligns task context with model expectations using prefill logits.
-  • NPU-optimized draft reuse (Section 3.4) maximizes NPU utilization and reduces verification costs by reusing plausible tokens from rejected drafts.
+To enable efficient RAG inference with end-to-end NPU offloading, sd.npu is designed as a runtime system that couples speculative decoding with NPU execution.
+Its key idea is a system–algorithm co-design: at the system level, sd.npu reduces graph-switching overhead, and at the algorithmic level, it introduces an NPU-optimized speculative decoding scheme.
+Architecture and workflow.
+Figure 8 illustrates the overall architecture.
+It intercepts the standard prefill-decoding pipeline and orchestrates speculative decoding workloads tailored for NPU execution with the following workflow:
+$\bullet$ Offline initialization.
+sd.npu partitions the LLM into several blocks, each containing a contiguous set of Transformer layers.
+A one‑time warm‑up profiling characterizes each block’s latency and memory footprint under different chunk sizes.
+This profile is used to determine the optimal partitioning and graph combination for the device’s NPU configuration.
+sd.npu also maintains a token database that includes chat records and relevant corpus to facilitate draft generation.
+$\bullet$ Online execution. When an augmented user request arrives, sd.npu performs the following stages:
+Prefill.
+The request is partitioned into fixed‑size chunks following the approach of prior mobile LLM systems (work:mllm; work:heterollm).
+NPU executes these chunks using a long‑sequence optimized (prefill) graph.
+Progressive graph scheduling4.2).
+During prefill, the system gradually transitions to the decoding‑optimized graph through block-by-block switching, overlapping I/O costs with prefill computation (§4.2).
+Distribution calibration.
+After prefill, the system collects the output logits and applies in‑context distribution calibration (§4.3) to augment the draft database.
+This procedure aims to reduce rejection-caused recomputation by aligning drafts’ token statistics with the model’s conditional distribution.
+Speculative decoding.
+With the draft database, sd.npu performs speculative decoding using the decoding‑optimized graph.
+It iteratively constructs tree-structured drafts and verifies them via a confidence-based policy that identifies and reuses valuable tokens (§4.4).
+When an end‑of‑sequence token is verified, sd.npu returns the merged output and commits the interaction into the database to facilitate future generation.
 4.2. Progressive Graph Scheduling
 Challenges of fixed-shape compute graphs.
-As discussed in Section 2.3, mobile NPUs require fixed-shape compute graphs, forcing the prefill input to be chunked for variable-length requests.
-This constraint also forces prefill and decoding to share the same graph despite their distinct input patterns: prefill processes long sequences (e.g., chunk size 256), while decoding handles short inputs (e.g., chunk size 32).
-Using a unified graph thus leads to inefficiency.
-To address this, we consider two specialized graphs: $G^{1}$ optimized for prefill and $G^{2}$ optimized for decoding.
+As discussed in §3.1, mobile NPUs require static compute graphs, forcing prefill and decoding to share the same graph despite their distinct input patterns: prefill processes long sequences (e.g., input length 256), while decoding handles short inputs (e.g., input length 1).
+Using a shared graph thus leads to inefficiency, while keeping two graphs resident in memory is wasting limited RAM budget.
+To address this, we consider switching between two specialized graphs: $G^{1}$ optimized for prefill and $G^{2}$ optimized for decoding.
 We denote the computation latency of a forward pass of model $\operatorname{P}$ on graph $G^{x}$ ($x\in\{1,2\}$) as $T_{\text{compute}}(\operatorname{P};G^{x})$.
 […]
-Switching graph progressively. sd.npu gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
-  • Graph equivalence by repetition: executing $G^{2}$ multiple times on shorter inputs yields the same result as one execution of $G^{1}$ on longer inputs, allowing early switching.
-As sd.npu maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
-  • Chunk-level scheduling: multiple block computation can be overlapped with a single block loading.
+Progressive switching. Instead, sd.npu amortizes the heavy switching cost across the entire prefill through two insights:
+$\bullet$ Graph equivalence: executing $G^{2}$ multiple times on shorter inputs is mathematically equivalent to a single execution of $G^{1}$ on longer inputs, allowing early switching.
+As sd.npu maximizes draft lengths (§4.4), $G^{2}$ will not be too short (e.g., chunk/4 of $G^{1}$), avoiding performance degradation.
+$\bullet$ Chunk decomposition: through fine-grained block partitioning and scheduling, the I/O payload of a single block can be masked by several chunked prefill iterations.
 (a) Progressive switching overlaps $G^{2}$ loading with computation across blocks.
 […]
-Figure 8. Progressive graph scheduling example with $N=6$ and two chunks. (a) shows the workflow of chunk 1 computation; (b) compares prefill pipelines under different scheduling schemes.
+Figure 9. Progressive graph scheduling example with $N=6$ and two chunks. (a) workflow of computing chunk 1; (b) pipelines comparison under different scheduling.
 Figure 9 illustrates progressive scheduling with $N=6$ and two input chunks.
@@
 The same procedure applies to chunk 2.
-In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency (Figure 8(b)).
+In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency, as shown in Figure 9(b).
 Problem Formulation.
-Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, sd.npu aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round.
+Formally, given $N$ model blocks (decided by offline profiling) that sequentially execute prefill for $C$ chunks, sd.npu aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round.
 The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
 Scheduling methodology.
-This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (Korf, 1998) (proofed in Appendix A).
+This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (work:partition).
 sd.npu addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute.
@@
 4.3. In-Context Distribution Calibration
-Challenges of language divergence.
-As discussed in Section 2.1, leveraging similarity on mobile LLMs is hindered by distribution divergence between task context and LLM modeling.
-Models often interpret the context with different lexical patterns, thus reducing the acceptance ratio of drafts directly generated from the context.
-These lexical gaps appear prominently at the token level.
-For example, in Figure 9, a user request states “… Clinton’s security detail has arrived at Des Moines …,” while the LLM, tasked with summarizing, prefers “Clinton arrived at Des Moines.” For classical retrieve-based speculative decoding approaches (Saxena, 2023; He et al., 2024; Hu et al., 2024), draft “Clinton’s security” is rejected due to lexical mismatch between token ’s and target arrived.
-This results in only one token is generated (arrived) by one forward pass as the following draft security is discarded, reducing decoding efficiency.
-Such mismatch problem arises in various mobile scenarios such as auto-reply or RAG-based QA, where human-written contexts diverge from the model’s distribution.
-Figure 9. Overview of in-context distribution calibration with a summarization example. :
-To tackle this issue, one may forcibly accept drafts with lexical gaps based on semantic-level similarity.
-However, setting a stringent similarity threshold results in a low acceptance ratio with only marginal performance gains, whereas a loose threshold causes non-negligible accuracy loss, e.g., drops from 0.9 to 0.75 if candidates are forcibly accepted ignoring the distribution divergence (Holsman et al., 2025; Bachmann et al., 2025).
-This necessitates calibration of draft and model distributions.
+Wasted computation due to rejection.
+To achieve lossless acceleration, speculative decoding rejects all draft tokens after the first incorrect token.
+Each rejection forces recomputation and falling back to token-by-token decoding, neutralizing speculative parallelism and wasting NPU cycles.
+This verification strategy requires lexical identity, resulting that semantically valid drafts are rejected.
+This lexical mismatches frequently arise in mobile RAG workloads such as summarization and QA, where retrieved text originates from human-written content with phrasing styles that deviate from the model’s learned patterns.
+As shown in Figure 10, given a draft “Clinton’s security detail has arrived at Des Moines,” the model’s own distribution predicts “Clinton arrived at Des Moines.” Although the semantics align, lexical mismatch at the token level (e.g., ’s vs.
+arrived) triggers rejection (work:pld; work:rest; work:sam).
+The verification then produces only one accepted token (arrived), forcing recomputation for the rest (security detail).
+Across NPU pipelines, this leads to redundant graph executions and degraded decoding throughput.
+Figure 10. Overview of in-context distribution calibration with a summarization example.
+Design goal.
+We aim to reduce wasted NPU computation caused by draft rejection, without compromising semantic accuracy.
+The key idea is to align drafts with the model’s intrinsic distribution such that they better match what the model is likely to verify.
 A straightforward solution is to precompute a calibrated database by running large volumes of user inputs through the LLM, which suffers from poor generalization to unseen queries and requires prohibitive storage and offline processing.
-Another method is to fine-tune the LLM to adapt to specific task scenarios, which also consumes lots of resources and energy and is hard to cover the wide diversity of mobile tasks.
-Leveraging prefill logits.
-To avoid heavy offline or fine-tuning costs, our key observation is that after prefill completes, next-token logits for each token in the entire context have already been calculated.
-These logits reflect how LLM encodes task semantics and can be used to approximate the model distribution.
-To that end, we can build a calibrated database for drafting with negligible performance overhead.
-Therefore, we propose in-context distribution calibration, an online, lightweight method that dynamically aligns context with the LLM distribution using prefill logits.
+Another method is to fine-tune the LLM to adapt to specific scenarios, which also consumes lots of resources and is hard to cover the wide diversity of RAG tasks.
+Leverage prefill logits.
+During prefill, NPUs already compute next-token logits for all context tokens as part of the model’s forward pass.
+These logits encode the LLM’s internal probability distribution and naturally reveal how the model interprets the current task context.
+By reusing these prefill logits, we can estimate the model’s local token transition space and calibrate the retrieval distribution without additional forward passes or fine-tuning.
+We call this mechanism in-context distribution calibration, an online, lightweight technique that exploits prefill logits to dynamically align drafts with model-side token probabilities, effectively avoiding wasteful computation on rejected drafts.
 Workflow.
 After prefill, sd.npu first collects predicted logits for all positions.
-Then sd.npu builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
+Then sd.npu builds a token-level calibration structure that approximates the model’s local generation distribution by 2-step depth-first search:
 […]
@@
 […]
-Overhead analysis.
-The main costs come from sampling and storing calibrated results.
-In our experiments, sampling incurs an amortized latency of about 2 ms per input token, while building and maintaining calibrated tokens peak below 500 MB memory for a 3B model (Section 4.3).
-3.4. NPU-Optimized Draft Reuse
-Challenges of short potentially acceptable drafts.
-Existing studies on R-SD focus on improving draft quality, proposing techniques such as pruning draft tree by estimating token confidence (Somasundaram et al., 2024; Hu et al., 2024; Wu et al., 2025a).
-However, as discussed in Section 2.3, a large proportion of drafts generated by existing methods have short lengths, which is harmful to the NPU efficiency.
-Naively lengthening drafts by loosing retrieving conditions cannot obtain performance benefits as irrelevant tokens will be rejected in verification.
-Reuse plausible drafts.
-As highlighted in Section 2.1, the model output usually shares similar semantics with the task context despite existing lexical gaps.
-Although these gaps cause rejection due to token-level misalignment, the rejected tokens include correct semantics and may be accepted in the next verification.
-sd.npu identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
-Table 2.
-Given draft $T,x_{1}\dots x_{n}$ which misaligns with target $z_{1}\dots z_{n+1}$ at position $e$, the possible patterns of segment $z_{e}\dots z_{n}$ and corresponding reusable tokens.
-In cases missing and synonymous, tokens after $x_{e+\epsilon}$ are omitted for simplicity.
+4.4. NPU-Optimized Draft Extension
+Short drafts underutilize NPU.
+NPUs are optimized for dense, long-running workloads that amortize weight loading and maximize matrix reuse across many tokens.
+However, existing workloads are typically short: most drafts contain 2-8 tokens, well below the length needed to saturate systolic arrays, as shown in Figure 5.
+Simply lengthening drafts is ineffective because irrelevant tokens are immediately rejected, forcing recomputation and wasting NPU cycles.
+Therefore, extending drafts with plausible tokens is crucial for efficient NPU execution.
+Observation: rejected tokens are not always incorrect.
+While SD rejects drafts with lexical mismatches, these tokens often encode semantically correct continuations.
+For example, when summarizing “Clinton’s security detail has arrived at Des Moines,” the model may reject “security” due to token-level mismatch, yet the phrase “Clinton arrived at” appears again in subsequent verification.
+Thus, the earlier rejected segment partially overlaps with future correct output, representing “wasted” computation that could be reused to extend future drafts and reduce underutilized NPU cycles.
 Case Pattern Reusable Ratio
@@
 Redundant $z_{e}\dots z_{n}$ $\emptyset$ 61.5%
-Rejection causes analysis.
-We first analyze the rejection reasons to explore how to identify plausible tokens.
-Let $\mathbf{x}=[T,x_{1},\dots,x_{n}]$ be the draft sequence where $T$ is the most recently accepted token, and $\mathbf{y}=[y_{1},\dots,y_{n+1}]$ be the model’s predictions.
-Suppose the expected sequence is $\mathbf{z}=[z_{1},\dots,z_{n+1}]$ and rejection occurs at position $e$ due to a mismatch: $x_{e}\neq y_{e}$.
-Thus, the accepted sequence is $x_{1},\dots,x_{e-1},y_{e}$.
-According to the patterns of the expected sequence $z_{e}\dots z_{n}$, we categorize reuse opportunities into three general cases and report their distribution, which is obtained by manually labeling the rejection causes from a user case of summarization.
-Table 2 summarizes the results, showing that nearly 40% of the cases include reused sequence portions.
-  • Missing: the expected sequence includes additional tokens $z_{e}\dots z_{e+\delta}\ (\delta\geq 0)$ not present in the draft, while the remaining draft tokens $x_{e}\dots x_{e+\epsilon}\ (\epsilon\geq 1)$ match the expectation.
-Thus, the expected sequence is $z_{e}\dots z_{e+\delta}x_{e}\dots x_{e+\epsilon}$.
-If the combined sequence $z_{e+\delta}x_{e}\dots x_{e+\epsilon}$ does not exist in the database, next retrieval will fail with no drafts generated.
-This could be recovered by reserving $x_{e}\dots x_{e+\epsilon}$.
-  • Synonym: the draft includes a semantically equivalent but lexically different segment $x_{e}\dots x_{e+\gamma-1}\ (\gamma\geq 1)$, replaced in the expected response by $z_{e}\dots z_{e+\delta}$.
-Similar failure occurs if the post-synonym tokens $x_{e+\gamma}\dots x_{e+\epsilon}$ cannot be retrieved from database using $z_{e+\delta}$.
-In this case, $x_{e+\gamma}\dots x_{e+\epsilon}$ should be preserved.
-  • Redundant: the entire draft $x_{e}\dots x_{n}$ is incorrect and fully diverges from the expected path. This is the only case where discarding the sequence is appropriate.
-Reuse methodology.
-To leverage reusable tokens in the first two cases, sd.npu aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
+Table 2.
+Given draft $T,x_{1}\dots x_{n}$ which misaligns with target $z_{1}\dots z_{n+1}$ at position $e$, the possible patterns of segment $z_{e}\dots z_{n}$ and corresponding reusable tokens.
+In cases missing and synonymous, tokens after $x_{e+\epsilon}$ are omitted for simplicity.
+Empirical analysis of reuse opportunities.
+We analyze rejection causes by comparing drafts $\mathbf{x}=[T,x_{1},\dots,x_{n}]$ with model predictions $\mathbf{y}=[y_{1},\dots,y_{n+1}]$ and expected sequences $\mathbf{z}=[z_{1},\dots,z_{n+1}]$.
+Suppose a rejection occurs at position $e$ where $x_{e}\neq y_{e}$.
+The accepted prefix is $x_{1},\dots,x_{e-1},y_{e}$.
+Our manual annotation across a summarization dataset (Table 2) reveals that nearly 40% of rejected tokens appear again within subsequent decoding steps, indicating that they can be safely reused to extend NPU workloads.
+Specifically, we categorizes three major rejection patterns in Table 2:
+$\bullet$ Missing: The expected output includes additional tokens absent from the draft, but subsequent draft tokens remain valid. Reusing $x_{e}\dots x_{e+\epsilon}$ preserves the valid suffix.
+$\bullet$ Synonym: The rejected tokens are semantically equivalent but lexically different (e.g., “car” vs.
+“vehicle”).
+Retaining the following segment $x_{e+\gamma}\dots x_{e+\epsilon}$ allows reusing once the synonym mismatch resolves.
+$\bullet$ Redundant: The entire segment diverges from the target sequence and must be discarded.
+Reuse strategy.
+sd.npu dynamically identifies reusable token segments from previously rejected drafts and appends them to new drafts to form NPU-friendly workloads.
+Formally, sd.npu aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
 Enumerating all possible reuse strategies is of $O(N^{3})$ complexity, which is not suitable for online scheduling.
@@
 Therefore, we identify and retain the longest segment of $\mathbf{x}$ such that $x_{i}=y_{i}$ for all $i\in[e+\gamma,e+\epsilon]$, by solving:
-(3) $\gamma,\epsilon=\underset{\gamma,\epsilon}{\text{argmax}}(\epsilon-\gamma)\quad\text{s.t. }y_{i}=x_{i},\ \forall\ i\in[e+\gamma,e+\epsilon]$
-Each reused segment is reserved for at least one iteration, and is discarded when the cumulative draft length of current decoding step exceeds a predefined threshold.
-By maximizing draft length with potentially acceptable tokens, this reuse mechanism directly enhances NPU utilization and reduces verification times.
-The overhead of sd.npu’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.
+(3) $\gamma,\epsilon=\underset{\gamma,\epsilon}{\text{argmax}}(\epsilon-\gamma)\quad\text{s.t. }y_{i}=x_{i},\ \forall\ i\in[e+\gamma,e+\epsilon].$
+This reused segment is retained across at least one decoding step and discarded once its accumulated lifetime or total draft length exceeds a predefined threshold.
+Thus, sd.npu effectively reuse wasted tokens and reduce underutilized ones, saving 70.5% of token computation without sacrificing accuracy.
+The additional overhead is discussed in § 5.4.
 5. Evaluation
-4.1. Implementation and Setup
-Implementation.
-We implement sd.npu  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code.
-We reuse the NPU backend of mllm, including neural kernels and memory managers, to perform prefill and draft verification.
-sd.npu inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities.
-Our modular design allows easy porting to other mobile LLM inference frameworks that face NPU inefficiencies during decoding.
-To enable NPU-optimized speculative decoding, we augment mllm with customized kernels and sampling strategies.
-In order to accelerate drafting, we implement a suffix automaton to identify potential suffix matches (Mohri et al., 2009).
-To support NPU graph switching, we integrate the graph saving and loading APIs provided by the Qualcomm QNN SDK (Qualcomm, 2025a).
+Figure 11. Per-token latency (ms/token) comparison on different datasets, devices and models.
+5.1. Experiment Setup
 Hardware.
@@
 Datasets.
-We evaluate four common mobile tasks collected from prior work: document summarization (summary), RAG-based question answering (rag), UI automation (ui) and automatic message reply (tweet(Xia et al., 2024; Xing et al., 2024; Salemi et al., 2023).
-Models. We use Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct and LLaMA3.2-3B-Instruct (Qwen et al., 2025; Grattafiori et al., 2024), exploring the effect of different model size.
+We evaluate four common mobile tasks collected from prior work: document summarization (summary(dataset:mt_bench), RAG-based question answering (rag(dataset:mt_bench), UI automation (ui(dataset:ui) and automatic message reply (tweet(dataset:lamp).
+Models. We use Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct (qwen25) and LLaMA3.2-3B-Instruct (llama3), exploring the effect of different model size.
 Baselines.
-We compare sd.npu against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation.
-sd.npu, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly.
-EdgeLLM (Xu et al., 2025a) optimizes model-based SD mainly to alleviate memory constraints.
-With sufficient memory, its performance reduces to that of EAGLE and does not address the memory-bound issue of decoding, thus making it an uninformative baseline for our settings.
+We compare with mllm and two on-device LLM inference systems that optimizes speculative decoding: SpecInfer (work:specinfer) for model-based SD and NG+ (work:ng+) for retrieval-based SD.
+To eliminate performance gaps from backend design differences, we implement their algorithms atop mllm  [^1] Other existing mobile LLM inference systems with NPU offloading support such as PowerInfer-v2 (work:powerinfer2) and HeteroLLM (work:heterollm) are not open-source, making them unsuitable for direct implementation.
+sd.npu, as a system-level optimization, is orthogonal and integrable with T-MAN (work:tman), so we do not compare it directly.
 .
-All baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (Xu et al., 2025b; Chen et al., 2025a; Li et al., 2024a; Xue et al., 2024).
-We categorize these baselines by whether they use CPU for decoding or apply speculative decoding:
-  • (1)NPU vanilla: use NPU for both prefill and decoding222mllm (Xu et al., 2025b) and other on-device LLM engines  (Chen et al., 2025a; Xue et al., 2024) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality.
+While all baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (work:mllm; work:heterollm; work:transformer_lite; work:powerinfer2), we evaluate decoding baselines on different backends, i.e., CPU and NPU to augment our experiments.
+In summary, our baselines include:
+(1) mllm(NPU): use NPU for both prefill and decoding222mllm (work:mllm) and other on-device LLM engines  (work:heterollm; work:powerinfer2) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality.
 sd.npu follows this pattern even in the NPU vanilla mode.;
-  • (2)CPU vanilla: use NPU for prefill and CPU for decoding, adopted by recent frameworks such as mllm and HeteroLLM (Xu et al., 2025b; Chen et al., 2025a);
-  • (3)NPU-SAM: apply retrieval-based speculative decoding (SAM (Hu et al., 2024)) to NPU vanilla.
-SAM drafts from both historical and local contexts, serving as a standard retrieval-based baseline without mobile-specific optimization;
-  • (4)NPU-SAM: similar to (3) but decoding on the CPU;
-  • (5)NPU-EAGLE: apply model-based speculative decoding (EAGLE-2 (Li et al., 2024b)) to NPU vanilla.
-EAGLE-2 is the SOTA model-based SD approach that uses both tokens and hidden states for drafting through a pretrained transformer-like model;
-  • (6)CPU-EAGLE: similar to (5) but decoding on the CPU.
-Since EAGLE-2 requires a separate draft model that is unavailable for selected LLMs, baselines (5) and (6) are simulated by estimating the expected speed via amortizing the forward costs of the target and draft model.
-We approximate the draft model as a single transformer layer with a head layer of the target LLM, assuming a fixed maximum acceptance ratio of 5.5, as reported in its paper (Li et al., 2024b).
+(2) mllm(CPU): use NPU for prefill and CPU for decoding;
+(3) NG+(NPU): apply retrieval-based SD from NG+ (work:ng+) to mllm-NPU. We use SAM (work:sam) as retrieval algorithm, which is the SOTA method to draft from both request prompt and extra database;
+(4) NG+(CPU): similar to (3) but decoding on CPU;
+(5) SpecInfer(NPU): apply model-based SD from SpecInfer (work:specinfer) to mllm-NPU.
+We use EAGLE-2 (work:eagle2) as draft model, which is the SOTA model-based SD approach that uses a pretrained transformer-like model for drafting;
+(6) SpecInfer(CPU): similar to (5) but decoding on CPU.
+Hyper-parameter settings.
+In practice, we observe that the $N=2$ partitioning is sufficient for our selected models and devices to fully overlap the graph switching costs by computation.
+Therefore, we fix the model partitioning schema of $N=2$ across all evaluation.
+For the retaining threshold in draft extension technique, we set a limit of 32, which is the least draft length to saturate NPU as shown in Figure 5.
 Metrics and Configuration.
 We evaluate average per-token latency, energy consumption and peak memory.
-We exclude accuracy comparison as sd.npu achieves lossless generation (Leviathan et al., 2023a).
+We exclude accuracy comparison as sd.npu follows the strict verify schema in SD (work:sd_is_lossless) and achieves lossless generation.
 Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms.
@@
 Per-Token Latency. sd.npu is faster and achieves improvements ranging 1.06–3.81×, as shown in Figure 11.
-Figure 10. Per-token latency (ms/token) comparison on different datasets, devices and models. :
-$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×.
+$\bullet$ Compared to baselines without SD (mllm(NPU) and mllm(CPU)), sd.npu consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×.
 These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs.
-Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is memory-bound which is preferred by SD.
+Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is less compute-intensive.
 Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
-$\bullet$ Compared to baselines with SD, sd.npu consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE.
-This arises because SAM ignores the ineffectiveness of computing short drafts on NPUs while EAGLE suffers from the latency of running another parametric model.
-Improvements of applying SD on CPU vanilla are moderate (0.92–1.13× for SAM, 0.56–1.18× for EAGLE), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
-Energy Consumption. sd.npu is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
-$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×.
+$\bullet$ Compared to baselines with SD, sd.npu consistently improves by 1.11–2.53× than NG+, 1.09–1.80× than SpecInfer.
+This arises because NG+ ignores the inefficiency of computing short drafts on NPUs while SpecInfer suffers from the latency of running another parametric model.
+Improvements of applying SD on CPU are moderate (0.92–1.13× for NG+, 0.56–1.18× for SpecInfer), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
+Energy. sd.npu is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 12).
+$\bullet$ Compared to baselines without SD (mllm(NPU) and mllm(CPU)), sd.npu significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×.
 The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs.
 On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
-$\bullet$ Compared to baselines with SD, sd.npu consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×).
-Energy saving of SAM is limited due to frequent verification incurred by rejected drafts.
-EAGLE suffers from additional computation required by drafter, amplifying energy overhead.
-For instance, for Qwen2.5 0.5B, sd.npu reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
+$\bullet$ Compared to baselines with SD, sd.npu consumes significantly less energy than NG+ (1.07–3.07×) and SpecInfer (1.77–4.71×).
+Energy saving of NG+ is limited due to frequent draft rejection.
+SpecInfer suffers from energy overhead amplified by drafter computation.
+For instance, for Qwen2.5 0.5B, sd.npu reduces energy 1.40–3.50× compared with mllm(NPU), while NG+ reduces energy by only 1.17–1.54× and SpecInfer even increases energy consumption.
 Comparison Across Devices.
 We find that NPU computation on Redmi K70 Pro is much slower than on other devices, while Redmi K60 Pro and Oneplus 13 demonstrate relatively similar performance patterns.
-This results in less profound latency improvements on Redmi K70 Pro (e.g., 1.14–2.91× compared to NPU vanilla and 1.13–1.64× compared to CPU vanilla) as compute becomes the bottleneck.
+This results in less profound latency improvements on Redmi K70 Pro (e.g., 1.14–2.91× compared to mllm(NPU) and 1.13–1.64× compared to mllm(CPU)) as compute becomes the bottleneck.
 In terms of energy consumption, Oneplus 13 consumes more energy than Redmi K60 Pro, e.g., average 53.85 J and 56.52 J per request for Qwen2.5 1.5B.
-Figure 11. Energy consumption (J) comparison on different datasets, devices and models. :
+Figure 12. Energy consumption (J) comparison on different datasets, devices and models.
 5.3. Memory Overhead
-Figure 12. Peak memory comparison (GB) on different devices and models. :
-As shown in Figure 12, sd.npu maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations.
-For example, sd.npu requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB.
+As shown in Figure 13, sd.npu maintains a memory footprint close to mllm(NPU) with an additional overhead less than 500 MB across all evaluations.
+For example, sd.npu requires 3.14 GB for Qwen2.5 1.5B, nearly identical to mllm(NPU) (3.05–3.12 GB), while mllm(CPU) requires over 6 GB.
 This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
 The memory overhead of sd.npu primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory.
-In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that sd.npu delivers superior efficiency without sacrificing memory scalability.
+In contrast, SpecInfer’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that sd.npu delivers superior efficiency without sacrificing memory scalability.
+Figure 13. Peak memory comparison (GB) on different devices and models.
 5.4. Latency Breakdown
 To analyze the runtime characteristics and latency overhead of sd.npu, we decompose the end-to-end inference latency into prefill, decoding and potential overhead.
-Measurements are conducted on Oneplus 13 under the summary benchmark, with results shown in Figure 13.
-Figure 13. Breakdown of inference latency (ms). :
-The results show that sd.npu significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead.
-Specifically, decoding of NPU vanilla baseline consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization.
-sd.npu directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
+Measurements are conducted on Oneplus 13 under the summary benchmark.
+Figure 14. Breakdown of inference latency (ms).
+The results (Figure 14) show that sd.npu significantly accelerates both prefill and decoding compared to mllm(NPU) baseline, with little additional overhead.
+Specifically, decoding of mllm(NPU) consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization.
+sd.npu directly addresses this by integrating NPU-optimized speculative decoding, reducing decoding latency by 5.25–8.15×.
 The prefill overhead is due to sampling tokens from logits for distribution calibration, which costs about 2 ms for each input token.
-The decoding overhead is caused by draft retrieval and tree construction.
+The decoding overhead is caused by draft tree construction.
 Benefited from the low time complexity of suffix automaton, our drafter costs less than 3 ms for each decoding step.
@@
 Effectiveness of three techniques.
-Experiments are conducted on Redmi K60 Pro using the summary benchmark across four baselines: (1) NPU vanilla (NPU), (2) NPU-SAM (+SD), (3) NPU-Graph (+G), and (4) NPU-SAM-Graph (+G+SD).
-Baselines (1)(2) are identical with those in Section 4.2, while (3)(4) integrate progressive graph scheduling into (1) and (2), respectively.
+Experiments are conducted on Redmi K60 Pro using the summary benchmark across four baselines: (1) mllm(NPU), (2) NPU-SD (+SD), (3) NPU-Graph (+G), and (4) NPU-SD-Graph (+G+SD).
+Baselines (2)(4) use SAM (work:sam) without our optimization, while (3)(4) integrate progressive graph scheduling into (1) and (2).
 Results are reported in Table 3.
-Table 3. Ablation experiments on average per-token latency (ms/token).
 Model NPU +G +SD +G+SD Ours
@@
 LLaMA3.2-3B 71.71 70.03 47.93 46.25 41.62
+Table 3. Ablation on average per-token latency (ms/token).
 We observe that progressive graph scheduling provides consistent speed improvements regardless of whether SD is applied.
-For example, for Qwen2.5-1.5B, graph scheduling reduces latency from 57.41 ms to 47.72 ms.
-The results also show that in-context distribution calibration and NPU-optimized draft reuse contribute to significant performance gains across three models, reducing latency from 25.35 ms to 15.08 ms for Qwen2.5-1.5B.
-Effectiveness in acceptance ratio.
-To analyze the effectiveness of techniques relevant to R-SD, we evaluate the acceptance ratio gains applying our techniques to existing methods (PLD (Saxena, 2023), SAM (Hu et al., 2024)).
-PLD drafts from local contexts only while SAM drafts from both local and historical contexts.
+Note that the performance gains are amortized across the end-to-end generation, thus its improvement is less pronounced for larger LLM whose computation is the main bottleneck, or small model that requires relatively short time to switch the graph compared to the whole generation latency.
+The results also show that our optimized SD contributes to significant performance gains across three models, reducing latency from 25.35 ms to 15.08 ms for Qwen2.5-1.5B.
+Effectiveness in accept length.
+To analyze the effectiveness of SD optimization, we evaluate the accept length gains applying our techniques to existing methods (PLD (work:pld), SAM (work:sam)).
+PLD drafts from current context (i.e., user request) only while SAM drafts from both current context and an extra database.
 Experiments are conducted on a cloud server equipped with A100 GPUs.
 The results in Table 4 show consistent improvements across benchmarks and models.
-Our techniques improve the acceptance ratio of PLD by 1.05–1.25×, improve SAM by 1.06–1.24×.
-These improvements directly translate to higher decoding efficiency, as a higher acceptance ratio reduces the number of tokens that must be recomputed.
-These results also demonstrate that in-context distribution calibration and draft reuse strategy are orthogonal enhancements with existing R-SD methods, making it effortless to integrate advanced techniques.
-Table 4. Ablation on acceptance ratio (average accepted tokens per verification).
+Our techniques improve the accept length of PLD by 1.05–1.25×, improve SAM by 1.06–1.24×.
+These improvements directly translate to higher NPU efficiency, as discussed in § 3.2.
+These results also demonstrate that our techniques provide orthogonal enhancements with existing SD methods, making it effortless to integrate advanced techniques.
 Dataset Method Q1.5B Q7B L1B
 summary PLD 2.53 1.55 1.42
-PLD+Ours 2.91 ($\times$1.15) 1.84 ($\times$1.18) 1.70 ($\times$1.19)
+Ours 2.91 (1.15×) 1.84 (1.18×) 1.70 (1.19×)
 SAM 3.60 1.96 1.68
-SAM+Ours 4.21 ($\times$1.17) 2.43 ($\times$1.24) 2.07 ($\times$1.24)
+Ours 4.21 (1.17×) 2.43 (1.24×) 2.07 (1.24×)
 rag PLD 3.19 1.63 1.77
-PLD+Ours 3.73 ($\times$1.17) 1.89 ($\times$1.16) 2.11 ($\times$1.19)
+Ours 3.73 (1.17×) 1.89 (1.16×) 2.11 (1.19×)
 SAM 4.42 2.29 2.25
-SAM+Ours 5.47 ($\times$1.24) 2.80 ($\times$1.22) 2.66 ($\times$1.19)
+Ours 5.47 (1.24×) 2.80 (1.22×) 2.66 (1.19×)
 ui PLD 2.28 1.64 1.57
-PLD+Ours 2.64 ($\times$1.16) 1.73 ($\times$1.05) 1.82 ($\times$1.16)
+Ours 2.64 (1.16×) 1.73 (1.05×) 1.82 (1.16×)
 SAM 3.56 2.44 1.72
-SAM+Ours 4.01 ($\times$1.13) 2.59 ($\times$1.06) 2.05 ($\times$1.20)
+Ours 4.01 (1.13×) 2.59 (1.06×) 2.05 (1.20×)
 tweet PLD 2.53 2.06 1.49
-PLD+Ours 3.17 ($\times$1.25) 2.52 ($\times$1.22) 1.72 ($\times$1.15)
+Ours 3.17 (1.25×) 2.52 (1.22×) 1.72 (1.15×)
 SAM 3.57 2.98 1.79
-SAM+Ours 3.82 ($\times$1.07) 3.70 ($\times$1.24) 2.18 ($\times$1.22)
+Ours 3.82 (1.07×) 3.70 (1.24×) 2.18 (1.22×)
 Q1.5B = Qwen2.5 1.5B, Q7B = Qwen2.5 7B, L1B = LLaMA3.2 1B.
+Table 4. Ablation on accept length.
 6. Related Work
-Mobile LLM inference optimization.
-The constrained resources of mobile devices require reducing computation to enable efficient on-device LLM inference (Xu et al., 2024; Subramanian et al., 2025; Ma et al., 2024; Thawakar et al., 2024; Xu et al., 2025a).
-Several approaches exploit different forms of sparsity.
-Mixture-of-experts leverages activation sparsity by computing only neurons important to the final output (Yi et al., 2025; Chen et al., 2025c; Wu et al., 2025b; Frantar and Alistarh, 2024).
-Early-exit reduces layer-level computation by skipping non-critical layers (Venkatesha et al., 2025; Fan et al., 2024).
-Model compression reduces model size by retaining only the most valuable layers (Yu et al., 2024; Smith et al., 2025).
-Token pruning eliminates redundant tokens to shorten the context (Liu et al., 2025b; Ma et al., 2023).
-sd.npu is compatible with these techniques, as it requires no architectural modifications to models or inputs.
-Mobile NPU offloading.
-Modern mobile SoCs are increasingly equipped with high-performance NPUs (OPPO, 2024; Samsung, 2024; Huawei, 2024), offering new opportunities for LLM acceleration (Xue et al., 2024; Chen et al., 2025a; Lu et al., 2025; Xu et al., 2025b; Yin et al., 2025; Wei et al., 2025).
-PowerInfer-V2 (Xue et al., 2024) integrates parameter swapping with NPU offloading to scale inference to larger LLMs.
-HeteroLLM (Chen et al., 2025a) introduces tensor partitioning strategies tailored to mobile GPUs and NPUs for efficient parallelization.
-While these studies primarily focus on accelerating prefill, sd.npu complements them by incorporating speculative decoding to optimize both prefill and decoding.
-Retrieval-augmented generation.
-Retrieval-augmented generation (RAG) enhances LLM’s knowledge by retrieving relevant documents from an external database, enabling effective context-aware generation (Lewis et al., 2020).
-Recent work explores advanced database architectures such as knowledge graph (Edge et al., 2024; Han et al., 2025; Zhang et al., 2025).
-Other efforts optimize retrieval with adaptive strategies, such as extracting fine-grained information and caching key documents based on user demands (Jeong et al., 2024; Seemakhupt et al., 2024; Liu et al., 2025a).
-RAG can be employed to augment the context for CAG.
-Speculative decoding uses LLM as a verifier to process multiple draft tokens in parallel, shifting memory-bound decoding to compute-bound (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023b).
-The main difference between these studies is how drafts are generated.
-Model-based methods employ an auxiliary parametric model, typically a smaller LLM (Li et al., 2024c, b; Cai et al., 2024; Du et al., 2024; Li et al., 2025b).
-Although these approaches improve the acceptance ratio, they introduce significant drafting overhead (Hu et al., 2024).
-Retrieval-based methods avoid this cost by using lightweight retrievers for drafting, but often suffer from low acceptance ratios (Saxena, 2023; Somasundaram et al., 2024; Hu et al., 2024; Chen et al., 2025b; Li et al., 2025a; Gritta et al., 2025).
-sd.npu advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
-6. Conclusions
-In this paper, we propose sd.npu, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation.
-sd.npu incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs.
-Extensive experiments demonstrate sd.npu achieves outstanding and consistent performance gains.
+LLM inference on mobile NPU.
+Mobile SoCs are increasingly equipped with high-performance NPUs (trend:oppo; trend:samsung; trend:huawei), offering new opportunities for LLM acceleration (work:powerinfer2; work:heterollm; work:bluelm; work:mllm; work:agent_xpu).
+mllm (work:mllm) and PowerInfer-V2 (work:powerinfer2) offload precision-insensitive operations to NPU, while HeteroLLM (work:heterollm) pipelines GPU and NPU.
+Test-time-compute NPU (work:ttc) improves response quality by NPU-aware quantization.
+T-MAN (work:tman) reduces quantization overheads by kernel refinement.
+sd.npu complements these work by incorporating system-algorithm co-design to optimize RAG inference.
+Retrieval-augmented generation enhances LLM’s knowledge using relevant documents from an external database, enabling effective context-aware generation (work:rag).
+Recent work explores advanced database architectures such as knowledge graph (work:graph_rag1; work:graph_rag2; work:graph_rag3).
+Other efforts optimize retrieval with adaptive strategies (work:adaptive_rag; work:edge_rag; work:webanns) and improve memory management for long contexts (work:infinigen; work:page_attention).
+Speculative decoding uses LLM as a verifier for parallel computing (work:early_speculative_decoding; work:speculative_sampling; work:speculative_decoding).
+Model-based SD uses an auxiliary parametric model for drafting, typically a smaller LLM (work:eagle; work:eagle2; work:medusa; work:glide; work:eagle_3).
+Although these approaches improve the accept ratio, they introduce significant drafting overhead (work:sam).
+Retrieval-based SD avoids this cost by using lightweight retrievers for drafting in the cost of low accept ratios (work:pld; work:pld+; work:sam; work:rapid; work:chunk_sd; work:dresd).
+sd.npu advances them through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
+7. Conclusion
+[…]
 Appendix A NP-hardness Proof of the Graph Scheduling Problem
-We prove that the graph scheduling problem described in Section 3.2 is NP-hard via a reduction from the 2-Partition problem (Korf, 1998).
+We prove that the graph scheduling problem described in Section 4.2 is NP-hard via a reduction from the 2-Partition problem (work:partition).
 Graph scheduling problem.